### v2

# Infinite Dimensional Topological-Holomorphic Symmetry in Three-Dimensions

Abstract.
We introduce a three-dimensional quantum field theory with an infinite-dimensional symmetry, realized explicitly through a centrally extended affine graded Lie algebra. This symmetry is a direct three-dimensional generalization of the chiral symmetry in the Wess-Zumino-Witten model. Upon performing radial quantization, we construct the Fock space of the theory and, via a three-dimensional analogue of the state–operator correspondence, we demonstrate that the algebra of local operators is endowed with the structure of a raviolo vertex algebra. Accordingly, this setup provides a framework for extending the methods of two-dimensional conformal field theory to three dimensions, and we expect it to lay the groundwork for exact methods in three-dimensional quantum field theory.
1. Introduction
Two-dimensional conformal field theories have become a foundational framework for both physics and mathematics. The modern formulation of conformal invariance in two dimensions began with the seminal work of Belavin, Polyakov, and Zamolodchikov [1]. Building on the representation theory of the Virasoro algebra, they introduced the concept of an algebra of local operators and used it to construct exactly solvable conformal theories, known as minimal models. This breakthrough sparked intense activity at the intersection of mathematical physics and statistical mechanics, as minimal models were soon identified with a variety of two-dimensional statistical systems at criticality.
In view of these developments, many efforts were devoted to extending the powerful tools of two-dimensional conformal field theory to higher dimensions. These attempts, however, encountered fundamental obstacles. The most immediate is Liouville’s rigidity theorem, which implies that for $d\geq 3$, all local conformal transformations extend to global ones, and the global conformal group is finite-dimensional. This rules out the existence of local infinite-dimensional conformal symmetry algebras in higher dimensions.
Given these obstructions, a prevailing strategy has been to isolate 2d chiral subalgebras within higher-dimensional theories, where infinite-dimensional symmetries can still be realized. A prominent example is provided by 44d $\mathcal{N}=2$ superconformal field theories, where a protected sector of local operators was shown in [2] to organize into a two-dimensional vertex operator algebra, enabling the use of chiral algebra techniques to obtain exact results. More recently, the celestial holography program has proposed that 44d scattering amplitudes can be reinterpreted as correlators in a 22d conformal field theory on the celestial sphere [3, 4], where the hope is that this perspective may allow the application of chiral algebra techniques to the problem of flat space holography [5, 6, 7, 8, 9].
The aim of this work is to revisit the pursuit of infinite-dimensional symmetry algebras in three-dimensional quantum field theories, but with a shift in perspective. Rather than looking for a subsector whose dynamics can be captured by a two-dimensional chiral algebra, the strategy is to generalize the very notion of a chiral algebra into an infinite-dimensional symmetry structure suited to describe three-dimensional dynamics. These algebras enjoy a generalized notion of chirality while remaining infinite-dimensional, and thus preserve many of the powerful features typically associated with chiral algebras.
The three-dimensional theories that we will consider are said to be partially topological and partially holomorphic. They first appeared in the context of twisted 3d $\mathcal{N}=2$ supersymmetric field theories [10] (see also [11, 12, 13, 14]) and more recently, in the contexts of Poisson vertex algebras [15, 16], factorization algebras [17] and twistorial field theories [18]. We will focus on a specific example introduced by the authors in [19], obtained via a localization procedure applied to holomorphic 2-Chern-Simons theory.
These theories are defined on three-dimensional manifolds equipped with a transverse holomorphic foliation (THF) [20, 21], meaning they are locally modeled on $\mathbb{R}$$\times{\mathbb{C}}$. At first glance, one might expect that the analytic power of holomorphicity could be recovered in this setting simply by requiring fields to be holomorphic along the $\mathbb{C}$} direction. However, this condition alone is insufficient to reproduce the infinite-dimensional structure that underpins the symmetry algebras of two-dimensional quantum field theories. To recover an analogue of this structure in three dimensions, one must impose a stronger constraint: a compatibility condition that ties holomorphicity to independence along the topological direction. But how exactly?
A clever way to approach this question was proposed in [22], where the authors observed that the infinite-dimensional nature of chiral algebras can be traced back to a simple cohomological fact [^1] See Appendix A for our conventions on cohomological notation.:
(1.0.1) $H^{(0,0)}_{\bar{\partial}}(\mathbb{C}\setminus\{0\})={\mathbb{C}}[z,z^{-1}]\,.$
In general, in the contex of quantum field theory, chiral algebras are generated by the modes of a holomorphic current $J$. The condition $\bar{\partial}J=0$ then implies that $J$ admits a Laurent expansion, precisely because the cohomology of $\bar{\partial}$ is isomorphic to ${\mathbb{C}}[z,z^{-1}]$. This observation suggests a natural strategy for exploring infinite-dimensional symmetry structures in the three-dimensional setting: namely, to ask whether the cohomology of a suitable differential on $\mathbb{R}$$\times{\mathbb{C}}$ gives rise to a space of series.
Taking coordinates $(\tau,z,\bar{z})$ on $\mathbb{R}$$\times{\mathbb{C}}$, Garner and Williams show in [22] that the degree-zero and degree-one cohomology of the differential
(1.0.2) $\mathrm{d}^{\prime}=\bar{\partial}+\mathrm{d}_{\tau}\,$,
each correspond to a certain space of series. This suggests that a three-dimensional field theory admitting a conserved current $J$ satisfying $\mathrm{d}^{\prime}J=0$ could, in principle, exhibit an infinite-dimensional symmetry algebra generated by the modes of $J$. While this observation provides a concrete path towards realising such an infinite-dimensional symmetry structure, [22] does not identify a concrete field theory realising such a current, leaving its realisation as an open question.
This is precisely where the three-dimensional topological-holomorphic field theory constructed in [19] enters the picture. It provides a concrete realisation of a quantum field theory with symmetry currents satisfying, schematically, a conservation law of the form $\mathrm{d}^{\prime}J=0$. In [19], it was already anticipated that the symmetries of the theory should form an infinite-dimensional algebra. What was missing, however, was a precise mathematical description of this structure. By recognizing that the differential constraint obeyed by the symmetry currents coincides with the one underlying the raviolo complex of [22], we show that this framework provides exactly the missing ingredient which allows for the explicit construction of the infinite-dimensional symmetry algebra.
More precisely, the condition $\mathrm{d}^{\prime}J=0$ allows one to expand the symmetry currents into formal series whose coefficients define an infinite collection of conserved charges, in direct analogy with the mode expansion of chiral currents in two-dimensions. In this context, one can also perform radial quantisation, thereby allowing the computation of commutators between the modes of the symmetry currents. The commutation relations are then shown to correspond to a centrally extended affine graded Lie algebra, which can be understood as the three-dimensional counterpart of the Kac–Moody algebra.
The existence of this algebra makes it possible to construct the Fock space of the theory by consistently selecting the creation operators among the modes of the symmetry currents. In the language of representation theory, this Fock space corresponds to the vacuum module of the algebra. As in the two-dimensional case, one can show that the states of this vacuum module are in one-to-one correspondence with local operators of the three-dimensional theory — the so-called state-operator correspondence. Altogether, these ingredients define a raviolo vertex algebra in the sense of [22], which can be regarded as a three-dimensional generalisation of an ordinary vertex algebra.
We expect that this explicit three-dimensional quantum field theory, whose local operators are organised by a raviolo vertex algebra, represents a concrete step toward studying a broader class of topological–holomorphic models in three dimensions. In close analogy with two-dimensional conformal field theories, where the vertex algebra structure provides control over the space of operators and, in some cases, enables exact results, the presence of a raviolo vertex algebra suggests that similar progress may be achievable in three dimensions. Although this construction is still at an early stage, it provides a concrete starting point for exploring exactly solvable three-dimensional quantum field theories.
Outline
The structure of the paper is as follows.
In Section 2, we review the Wess–Zumino–Witten (WZW) model and show how the symmetries of this theory give rise, through radial quantisation, to two copies of the Kac–Moody algebra. In Section 3, we introduce the three-dimensional topological-holomorphic field theory constructed in [19], and describe its symmetry structure. In Section 4, we present the raviolo formalism developed in [22] and adapt it to the three-dimensional setting relevant to our theory. In Section 5, we perform the radial quantisation of the three-dimensional theory and compute the commutation relations between the modes of the symmetry currents. In Section 6, we show that these commutation relations define a centrally extended affine graded Lie algebra, which we construct in detail. Finally, in Section 7 we construct the raviolo vertex algebra corresponding to the three-dimensional theory.
2. WZW model and the Kac-Moody Algebra
Given a compact, simple Lie group $G$ and a two-dimensional Riemann surface $\Sigma$, the Wess-Zumino-Witten model describes a sigma model with target space $G$, with an additional topological term. The action is given by [23]
[…]
where $g:\Sigma\rightarrow G$ is a smooth field, and $\tilde{g}:B\rightarrow G$ is an extension to a three-manifold $B$ with boundary $\partial B=\Sigma$. The integer $k\in\mathbb{Z}$ is known as the level of the theory. The first term is the standard sigma model kinetic term, while the second—known as the Wess–Zumino term—depends only on the homotopy class of the extension $\tilde{g}$.
A remarkable feature of this model is its invariance under the infinitesimal transformations
(2.0.2) $g(z,\bar{z})\mapsto g(z,\bar{z})+g(z,\bar{z}){\tilde{\alpha}}(z)-\alpha(\bar{z})g(z,\bar{z})\,$,
where $(z,\bar{z})$ are local complex coordinates on $\Sigma$, $\alpha(\bar{z})$ is a $\mathfrak{g}$-valued anti-holomorphic function and, ${\tilde{\alpha}}(z)$ is a $\mathfrak{g}$-valued holomorphic function. The presence of these symmetries imply the existence of conserved currents, which are given by
(2.0.3) $J(z)=kg^{-1}\partial_{z}g\,,\quad\bar{J}(\bar{z})=k\partial_{\bar{z}}gg^{-1}\,.$
The conservation of these currents follows from the equations of motion. Indeed, under the infinitesimal variation $g\mapsto g+g\tilde{\alpha}$ with arbitrary $\tilde{\alpha}$, one obtains the equation
(2.0.4) $\partial_{\bar{z}}(g^{-1}\partial_{z}g)=0\,$,
while the variation $g\mapsto g-\alpha g$ yields the complementary equation
(2.0.5) $\partial_{z}(\partial_{\bar{z}}gg^{-1})=0\,.$
Since $J(z)=kg^{-1}\partial_{z}g$ is holomorphic and $\bar{J}(\bar{z})=\partial_{\bar{z}}gg^{-1}$ is anti-holomorphic, both currents admit Laurent expansions on ${\mathbb{C}}^{\times}={\mathbb{C}}\setminus\{0\}$ of the form
(2.0.6) $J(z)=\sum_{n\in\mathbb{Z}}J_{n}z^{-n-1}\,,\quad\bar{J}(\bar{z})=\sum_{n\in\mathbb{Z}}\bar{J}_{n}\bar{z}^{-n-1}\,.$
In his seminal work [23], Witten showed that the modes $J_{n},\bar{J}_{n}$ of the currents generate a Kac-Moody algebra at level $k$. We now review this result explicitly for the holomorphic current $J(z)$; the same analysis applies to $\bar{J}(\bar{z})$. Our discussion follows closely the presentation of [24].
2.1. The Kac-Moody Algebra
The first step in deriving the commutation relations of the modes $J_{n}$ is to consider the Ward identity associated with the symmetry generated by the holomorphic current $J(z)$. This identity specifies how correlation functions transform under symmetry variations. Concretely, if $X$ denotes a product of fields and $\langle X\rangle$ their correlation function, the Ward identity reads
(2.1.1) $\delta\langle X\rangle=\int\langle(\delta s)\,X\rangle\,$,
where $\delta s$ is the variation density of the action. For the current $J(z)$, the relevant transformation is the infinitesimal right action
(2.1.2) $\delta_{\tilde{\alpha}}g(z,\bar{z})=g(z,\bar{z}){\tilde{\alpha}}(z)\,$,
with ${\tilde{\alpha}}(z)$ a $\mathfrak{g}$–valued holomorphic function of compact support on a disk $D\subset\Sigma$, vanishing outside. Under this transformation the action varies as
(2.1.3) $\delta_{{\tilde{\alpha}}}S=\frac{1}{4\pi i}\oint_{\partial D}\mathrm{d}z\operatorname{Tr}\big({\tilde{\alpha}}(z)J(z)\big),$
where we have used Stokes’ theorem (counterclockwise orientation). Next, since $J(z)=kg^{-1}\partial_{z}g$ is $\mathfrak{g}$–valued, we fix a basis $\{t_{a}\}$ with $\operatorname{Tr}(t_{a}t_{b})=2\delta_{ab}$ and $t_{a},t_{b}]=f_{ab}^{c}t_{c}$, and write
(2.1.4) $J(z)=\sum_{a=1}^{\dim\mathfrak{g}}J^{a}(z)t_{a}\,.$
With these conventions, (2.1.1) takes the form
(2.1.5) $\delta_{\tilde{\alpha}}\langle X\rangle=\frac{1}{2\pi i}\oint_{\partial D}\mathrm{d}z\,{\tilde{\alpha}}_{a}(z)\langle J^{a}(z)X\rangle\,.$
This relation may be used to determine the operator product expansion (OPE) of the currents. Indeed, under the right–acting variation (2.1.2), the current transforms as
(2.1.6) $\delta_{{\tilde{\alpha}}}J(z)=[J(z),{\tilde{\alpha}}(z)]+k\partial_{z}{\tilde{\alpha}}(z)\,.$
Hence, taking the product of fields to be $X=J^{b}(w)X^{\prime}$, with $X^{\prime}=\prod_{i}\mathcal{O}_{i}(u_{i})$ where $w\in D$ and all $u_{i}\notin D$, and where the $\mathcal{O}_{i}$ denote arbitrary operator insertions, the holomorphy of $\tilde{\alpha}$} on $D$ (extended by zero outside) ensures that $\delta_{\tilde{\alpha}}$ acts only on $J^{b}(w)$. The Ward identity then yields
(2.1.7) $\big\langle\big([J,{\tilde{\alpha}}](w)^{b}+k\,\partial{\tilde{\alpha}}^{b}(w)\big)X^{\prime}\big\rangle=\frac{1}{2\pi i}\oint_{\partial D}\mathrm{d}z\,{\tilde{\alpha}}_{a}(z)\big\langle J^{a}(z)J^{b}(w)X^{\prime}\big\rangle\,$,
from which one extracts the operator product expansion
(2.1.8) $J^{a}(z)J^{b}(w)=\frac{k\delta^{ab}}{(z-w)^{2}}+\frac{f^{ab}_{c}J^{c}(w)}{z-w}+\mathrm{reg}\,$,
where $\mathrm{reg}$ denotes terms regular as $z\to w$. This operator identity is valid inside any correlator, provided all other insertions lie outside the integration contour.
The OPE provides the means to compute commutation relations, and this is most naturally formulated within the framework of radial quantisation. Recall that in any quantisation scheme one chooses a time direction and associates to each spatial slice a Hilbert space of states. In Euclidean signature no preferred time direction exists, so that one may choose the radial coordinate to be the time variable [^2] In Minkowski spacetime one may work on the cylinder, where a conformal map to the plane identifies the cylinder time coordinate with the radial coordinate.. In this setting, states are defined on circles of fixed radius, which play the role of equal–time slices. Correlation functions are then identified as vacuum expectation values of radially ordered products of operators:
(2.1.9) $\langle a(z)b(w)\rangle=\langle 0|\mathcal{R}(a(z)b(w))|0\rangle\,$,
where the radial-ordering operator $\mathcal{R}$ is defined by
(2.1.10) $\mathcal{R}(a(z)b(w))=\begin{cases}\begin{aligned} &a(z)b(w)\quad\text{if}\quad|z|>|w|\\ &b(w)a(z)\quad\text{if}\quad|w|>|z|\end{aligned}\end{cases}\,.$
With this identification, we can relate the OPE with commutators. To this end, we consider the contour integral
(2.1.11) $\oint_{S_{w}^{1}}\mathrm{d}z\,\mathcal{R}(a(z)b(w))\,$,
where the contour is a small counterclockwise circle around $w$. Next, we split the integration contour into two concentric, fixed–radius circles: an outer one $C_{1}$ with $|z|>|w|$ and an inner one $C_{2}$ with $|z|<|w|$ traversed in the same direction. Radial ordering then implies
(2.1.12) $\begin{split}\oint_{w}\mathrm{d}z\,\mathcal{R}(a(z)b(w))&=\oint_{C_{1}}\mathrm{d}z\,a(z)b(w)-\oint_{C_{2}}\mathrm{d}z\,b(w)a(z)\\ &=[A,b(w)]\end{split}$
where we have defined the operator
(2.1.13) $A=\oint\mathrm{d}z\,a(z)\,.$
If we want this relation to hold as an operator identity, we must allow for an arbitrary number of additional fields to appear alongside $a(z)$ and $b(w)$ in a general correlator. In particular, the decomposition of the contour into $c^{1}$ and $C_{2}$ is justified only when $b(w)$ is the unique insertion with a singular OPE with $a(z)$ lying between the two contours.
By integrating equation (2.1.12), we obtain the commutator between two operators, each defined as the contour integral of a holomorphic field:
(2.1.14) $A,B]=\oint_{S_{0}^{1}}\mathrm{d}w\oint_{S_{w}^{1}}\mathrm{d}z\,\mathcal{R}(a(z)b(w))\,.$
From the above relation, one can compute the commutation relations for the modes in the Laurent expansion of the currents. The result is summarized in the following proposition:
2.1.1 Proposition.
The modes $J^{a}_{n}$ of the current $J(z)$ satisfy the algebra
(2.1.15) $J^{a}_{n},J^{b}_{m}]=f^{ab}{}_{c}\,J^{c}_{n+m}+kn\,\delta^{ab}\delta_{n+m,0}\,$,
which are precisely the commutation relations of the Kac-Moody algebra at level $k$.
Proof.
Writing the modes in terms of the current as
(2.1.16) $J_{n}^{a}=\frac{1}{2\pi i}\oint_{S^{1}}\mathrm{d}z\,z^{n}J^{a}(z)\,$,
we may express the commutator as in equation (2.1.14)
(2.1.17) $\begin{split}[J_{n}^{a},J_{m}^{b}]&=\frac{1}{(2\pi i)^{2}}\oint_{S_{0}^{1}}\mathrm{d}w\,w^{n}\oint_{S_{w}^{1}}\mathrm{d}z\,z^{m}\mathcal{R}(J^{a}(z)J^{b}(w))\\ &=\frac{1}{(2\pi i)^{2}}\oint_{S_{0}^{1}}\mathrm{d}w\,w^{n}\oint_{S_{w}^{1}}\mathrm{d}z\,z^{m}\left(\frac{k\delta^{ab}}{(z-w)^{2}}+\frac{f^{ab}_{c}J^{c}(w)}{z-w}+\mathrm{reg}\right)\,.\end{split}$
In the second line we replaced the radially ordered product by the OPE, as the singular part is invariant under radial ordering due to the antisymmetry of $f^{ab}_{c}$, and only this part contributes to the contour integral. Expanding $z^{m}$ around $w$ gives
(2.1.18) $z^{m}=\sum_{k=0}^{m}\binom{m}{k}\,w^{m-k}(z-w)^{k}\,$,
so that only the simple and double pole terms contribute to the residue. Using Cauchy’s theorem we then obtain
(2.1.19) $\begin{split}[J_{n}^{a},J_{m}^{b}]&=\frac{1}{2\pi i}\oint\mathrm{d}w\,w^{n}\left(k\,n\,\delta^{ab}\,w^{m-1}+f^{ab}_{c}J^{c}(w)\,w^{m}\right)\\ &=k\,n\,\delta^{ab}\,\delta_{m+n,0}+f^{ab}_{c}J_{m+n}^{c}\,.\end{split}$
This is precisely the Kac–Moody algebra at level $k$.
The fact that the modes of the chiral currents $J$ and $\bar{J}$ satisfy a Kac–Moody algebra has a number of significant consequences; among them is the fact that it makes the WZW model exactly solvable at the quantum level. Indeed, the chiral currents $J$ and $\bar{J}$ allow for the construction of the stress-energy tensor via the Sugawara construction [25, 26]. This tensor is conserved and traceless as an operator, ensuring full quantum conformal invariance. The combination of affine Kac–Moody and Virasoro symmetries imposes strong constraints on the theory. In particular, the Ward identities associated with the Kac–Moody symmetry lead to the Knizhnik–Zamolodchikov equation, which governs the correlation functions of primary fields [27]. In many cases, this equation can be solved exactly, yielding explicit expressions for these correlators — these are the conformal blocks. Since all other fields in the theory can be written in terms of primaries, this gives a complete solution for the WZW model.
2.2. 4d Chern-Simons Theory
The WZW model is closely related to 3d Chern Simons theory. This connection was first pointed out by Witten in [28], and subsequently made explicit in [29], where it was shown that a chiral WZW model arises naturally from Chern–Simons theory defined on a three-manifold with boundary.
Notably, the full (non-chiral) WZW model can also be obtained from four-dimensional Chern–Simons theory when the meromorphic 1-form is chosen such that the action takes the form [30]
(2.2.1) $S_{\mathrm{CS_{4}}}(A)=\int_{\Sigma\times\mathbb{C}P^{1}}\frac{{\mathrm{d}}z}{z}\wedge\operatorname{Tr}\big(A\wedge({\mathrm{d}}A+\tfrac{2}{3}A\wedge A)\big)\,$,
where $\Sigma$ is a two-dimensional manifold and $z$ is a coordinate on $\mathbb{C}P^{1}$. For an introduction to four-dimensional Chern–Simons theory, the reader may consult [31, 32, 33].
What makes this construction of the WZW model particularly interesting is that it provides a natural origin for the left- and right-acting symmetries described in equation (2.0.2). Indeed, the gauge symmetries of the gauge field $A$ at the four-dimensional level become, after localisation to $\mathbb{R}^{2}$, the holomorphic and anti-holomorphic symmetries of the two-dimensional field $g$. For a detailed explanation of this mechanism, we refer the reader to §5.1 of [19]. As we shall discuss in the next section, a similar analysis will allow us to identify the symmetries of the three-dimensional topological-holomorphic field theory.
3. 3d Topological-Holomorphic Theory
In this section, we introduce the main features of the three-dimensional theory constructed in [19], following the same analysis than in the previous section. Our goal is to show that the raviolo formalism is not merely a convenient tool, but in fact the natural framework for understanding the infinite dimensional symmetry structure of this theory, and more generally, of topological holomorphic field theories in three-dimensions.
Let $\mathfrak{G}=(\mathfrak{g}\xrightarrow{\mu_{1}}\mathfrak{h},\mu_{2})$ denote a Lie 2-algebra and let $\mathbb{G}=(\mathsf{H}\xrightarrow{\mathsf{t}}G,\rhd)$ the corresponding Lie 2-group. We consider a three-dimensional manifold $M$, which we take for now to be $M=\mathbb{R}\times\mathbb{C}$. The three-dimensional action is given by [19]
(3.0.1) $S[g,\Theta]=\int_{M}\operatorname{vol}_{M}\Big(\langle\partial_{z}(\partial_{\bar{z}}gg^{-1}),g\rhd\Theta_{\tau}\rangle-\langle\partial_{z}(\partial_{\tau}gg^{-1}),g\rhd\Theta_{\bar{z}}\rangle\\ +\langle\mu_{1}\big(\partial_{z}(g\rhd\Theta_{\bar{z}})\big),g\rhd\Theta_{\tau}\rangle\Big)\,$,
where $g\in C^{\infty}(M,G)$ is a $G$-valued smooth field, $\Theta\in\Omega^{1}(M,\mathfrak{h})$ an $\mathfrak{h}$-valued $1$-form, and $\langle\cdot,\cdot\rangle:\mathfrak{g}\times\mathfrak{h}\to\mathbb{R}$ a degree 1-1, non degenerate, invariant bilinear form. For an introduction to Lie $2$-groups and Lie $2$-algebras, we refer the reader to [34, 35, 36].
The appearance of the 1-form field $\Theta$, not as a gauge connection but rather as a matter field playing a role analogous to that of $g$, is closely tied to the higher categorical nature of the action. This reflects the categorical ladder = dimensional ladder principle [37, 38], which underlies the construction of the three-dimensional action (3.0.1). As we shall see, the presence of this higher-form field, together with the graded structure provided by the Lie 2-algebra, integrates seamlessly with the raviolo formalism developed in [22].
In complete analogy with the WZW model, the action (3.0.1) is invariant under both left- and right-acting symmetries. A direct computation shows that it is invariant under the transformations
(3.0.2) $g\mapsto g(1+\tilde{\alpha})\,,\quad\Theta\mapsto\Theta-\tilde{\alpha}\rhd\Theta+\tilde{\Gamma}\,$,
as long as the infinitesimal parameters $\tilde{\alpha}\in C^{\infty}(M,\mathfrak{g})$ and $\tilde{\Gamma}\in\Omega^{1}(M,\mathfrak{h})$ satisfy the constraints
(3.0.3) $\partial_{\bar{z}}\tilde{\alpha}+\mu_{1}(\tilde{\Gamma}_{\bar{z}})=0\,,\quad\partial_{\tau}\tilde{\alpha}+\mu_{1}(\tilde{\Gamma}_{\tau})=0\,,\quad\partial_{\tau}\tilde{\Gamma}_{\bar{z}}-\partial_{\bar{z}}\tilde{\Gamma}_{\tau}=0\,.$
These define the right-acting symmetries of the model. The left-acting symmetries take the form
(3.0.4) $g\mapsto(1-\alpha)g\,,\quad\Theta\mapsto\Theta-g^{-1}(1+\alpha)\rhd\Gamma\,$,
where the infinitesimal parameters $\alpha\in C^{\infty}(M,\mathfrak{g})$ and $\Gamma\in\Omega^{1}(M,\mathfrak{h})$ must satisfy
(3.0.5) $\partial_{z}\alpha=0\,,\quad\partial_{z}\Gamma_{\tau}=0\,,\quad\partial_{z}\Gamma_{\bar{z}}=0\,.$
Let us pause to highlight a few important observations. First, just like the fields in the theory, the symmetry transformations also come in pairs, once again reflecting the underlying categorical structure. Second, the left-acting symmetry pair is anti-holomorphic, exactly as in the WZW model. In contrast, the right-acting symmetry satisfies a more intricate condition: it involves derivatives with respect to $\bar{z}$ and $\tau$, along with a nontrivial appearance of the map $\mu_{1}$. This marks an important difference from the two-dimensional case. In particular, since the spacetime is now odd-dimensional, there is no natural chiral–anti-chiral decomposition, and we must carefully identify which of these symmetries plays the role of the chiral symmetry familiar from two dimensions—that is, which one gives rise to the infinite-dimensional symmetry algebra.
[…]
(3.0.6) $\partial_{\bar{z}}\tilde{L}_{z}-\mu_{1}\tilde{H}_{z{\bar{z}}}=0,\qquad\partial_{\tau}\tilde{L}_{z}-\mu_{1}\tilde{H}_{z\tau}=0,\qquad\partial_{\tau}H_{z\bar{z}}+\partial_{\bar{z}}H_{z\tau}=0\,.$
where we have introduced the currents
(3.0.7) $\tilde{L}_{z}=g^{-1}\partial_{z}g\,,\quad\tilde{H}_{z{\bar{z}}}=\partial_{z}\Theta_{\bar{z}}+\mu_{2}({\tilde{L}}_{z},\Theta_{\bar{z}}),\quad\tilde{H}_{z\tau}=\partial_{z}\Theta_{\tau}+\mu_{2}({\tilde{L}}_{z},\Theta_{\tau})\,.$
In a completely analogous manner, we can compute the currents associated to the left symmetry, given by
(3.0.8) $L_{\bar{z}}=-\partial_{\bar{z}}gg^{-1}-\mu_{1}(g\rhd\Theta_{\bar{z}})\,,\quad L_{\tau}=-\partial_{\tau}gg^{-1}-\mu_{1}(g\rhd\Theta_{\tau})\,$,
$\qquad H_{{\bar{z}}\tau}=g\rhd(\partial_{\bar{z}}\Theta_{\tau}-\partial_{\tau}\Theta_{\bar{z}}-[\Theta_{\bar{z}},\Theta_{\tau}])\,$,
which satisfy the conservation equations
(3.0.9) $\partial_{z}L_{\bar{z}}=0\,,\quad\partial_{z}L_{\tau}=0\,,\quad\partial_{z}H_{\bar{z}\tau}=0\,.$
3.1. Higher-currents as differential forms
The expressions we have obtained suggest a natural splitting of coordinates, treating $(\bar{z},\tau)$ and $z$ separately. More precisely, we can introduce the differential
(3.1.1) $\mathrm{d}^{\prime}=\bar{\partial}+\mathrm{d}_{\tau}\,$,
and decompose differential forms into components along $\mathrm{d}^{\prime}$ and along $\partial$. With this in mind, the currents associated with the right-acting symmetry can be written as
(3.1.2) $\tilde{L}=g^{-1}\partial g\,,\quad\tilde{H}=\partial\Theta+\mu_{2}(g^{-1}\partial g,\Theta)\,$,
and those associated with the left-acting symmetry as
(3.1.3) $L=-\mathrm{d}^{\prime}g\,g^{-1}-\mu_{1}(\Theta)\,,\quad H=g\rhd\left(\mathrm{d}^{\prime}\Theta-\tfrac{1}{2}[\Theta,\Theta]\right)\,.$
In terms of this decomposition, the conservation equation (3.0.6) takes the form
(3.1.4) $\mathrm{d}^{\prime}\tilde{L}-\mu_{1}(\tilde{H})=0\,,\quad\mathrm{d}^{\prime}\tilde{H}=0\,$,
and the differential constraints (3.0.5) satisfied by the symmetry parameters can be written as
(3.1.5) ${\mathrm{d}}^{\prime}\tilde{\alpha}+\mu_{1}(\tilde{\Gamma})=0\,,\quad{\mathrm{d}}^{\prime}\tilde{\Gamma}=0\,.$
Similarly, the conservation equation (3.0.9) can be written using the remaining differential as
(3.1.6) $\partial L=0\,,\quad\partial H=0\,$,
and the differential constraints (3.0.5) satisfied by the symmetry transformation parameters take the form
(3.1.7) $\partial\alpha=0\,,\quad\partial\Gamma=0\,.$
Let us note that this runs in complete parallel to what happens in the two dimensional case. Indeed, there, one can use the Dolbeault decomposition $\mathrm{d}$$\partial$$\bar{\partial}$ to define
(3.1.8) $J=kg^{-1}\partial g\,,\quad\bar{J}=k\bar{\partial}gg^{-1}$
such that the conservation equations become
(3.1.9) $\bar{\partial}J=0\,,\quad\partial\bar{J}=0\,$,
and the symmetry transformation parameters satisfy
(3.1.10) $\bar{\partial}\tilde{\alpha}=0\,,\quad\partial\alpha=0\,.$
At first glance, one might think this is just a convenient way to package our expressions. However, there is more to it: the existence of Laurent expansions for both the symmetry transformation parameter $\tilde{\alpha}(z)$ and the holomorphic current J(z)J(z)—which was crucial for realizing the symmetry algebra of the WZW model as the Kac-Moody algebra in Proposition 2.1.1—can instead be understood from a cohomological perspective.
Indeed, the condition $\bar{\partial}\tilde{\alpha}=0$ implies that [^3] See Appendix A for our conventions on cohomological notation.
(3.1.11) $\tilde{\alpha}\in H^{(0,0)}_{\bar{\partial}}((\mathbb{C}\setminus\{0\})\otimes\mathfrak{g})={\mathbb{C}}[z,z^{-1}]\otimes\mathfrak{g}\,\quad\Longrightarrow\quad\tilde{\alpha}=\sum_{n\in\mathbb{Z}}\tilde{\alpha}_{n}z^{n}\,.$
Furthermore, using the linear isomorphism [^4] We will make frequent use of this linear isomorphism. In two dimensions it amounts to writing $J=kg^{-1}\partial g$ as a one–form rather than as a $\mathfrak{g}$–valued function, a purely notational choice that eases the transition to the three–dimensional setting. Nevertheless, even in 3d it will at times be convenient to factor out the differential and regard these objects explicitly as either functions or forms.
(3.1.12) ${\mathrm{d}}z\wedge-:H^{(0,0)}_{\bar{\partial}}((\mathbb{C}\setminus\{0\})\otimes\mathfrak{g})\to H^{(1,0)}_{\bar{\partial}}((\mathbb{C}\setminus\{0\})\otimes\mathfrak{g})\,$,
one can also write a Laurent expansion for the current $J$, since $\bar{\partial}J=0$ implies
(3.1.13) $J\in H^{(1,0)}_{\bar{\partial}}((\mathbb{C}\setminus\{0\})\otimes\mathfrak{g})\,.$
In other words, the mode expansions of the current and symmetry parameters, can be understood as arising from their realization as nontrivial Dolbeault cohomology classes.
[…]
Crucially, to study the cohomology of $\mathbb{R}\times\mathbb{C}\setminus\{0\}$ using the splitting along $(\bar{z},\tau)$ and $z$, we must understand the decomposition of the entire complex $\Omega^{\bullet}(\mathbb{R}\times\mathbb{C}\setminus\{0\})$. While the Dolbeault decomposition for $\Omega^{\bullet}(\mathbb{C}\setminus\{0\})$ is a well established tool in complex geometry, the study of $\Omega^{\bullet}(\mathbb{R}\times\mathbb{C}\setminus\{0\})$ using the splitting defined by the differentials $\mathrm{d}^{\prime}$ and $\partial$ lies at the core of the raviolo formalism [22]. This is precisely why the raviolo framework provides the appropriate structure for analyzing our theory, as we shall discuss in detail in §4.
3.2. Transverse Holomorphic Foliation
In general, the theory (3.0.1) can be defined on any oriented 3-manifold $M$ equipped with local coordinate charts of the form $(\tau,z,\bar{z})\in\mathbb{R}\times\mathbb{C}$, such that the transition functions are given by
$(\tau,z,\bar{z})\mapsto(\tau^{\prime}(\tau,z,\bar{z}),z(z),\bar{z}(\bar{z}))\,.$
Such 3-manifolds are said to be equipped with a transverse holomorphic foliation $\mathcal{F}$. More precisely, we have
Definition 3.2.1.
Let $M$ denote a smooth manifold. A transverse holomorphic foliation (THF) on $M$ is a smooth foliation $\mathcal{F}$ of even codimension, whose leaves are equipped with a structure of a complex manifold. In other words, the foliation atlas of $(M,\mathcal{F})$ has equipped transition functions which are biholomorphic on the leaves. Vector fields in the tangent bundle $T\mathcal{F}\subset TM$ along the leaves are called the topological directions.
A field theory is said to be topological-holomorphic if its fields are defined on a smooth manifold $M$ equipped with a THF structure $\mathcal{F}$. It is clear that the kinematical and dynamical properties of a topological-holomorphic field theory depend crucially on the underlying THF structure. As we shall see in the upcoming section, the raviolo construction is done for general 3-manifolds which admit a THF.
4. The Raviolo
In this section, we review the raviolo formalism developed in [22]. As previously noted, this framework can be understood as the study of the cohomology of $\mathbb{R}\times\mathbb{C}\setminus\{0\}$, using the natural splitting of coordinates into $(\bar{z},\tau)$ and $z$. The core idea behind the raviolo formalism is that it provides a direct mechanism for constructing higher-dimensional vertex algebras (see also [39])—an algebraic structure that has proven remarkably powerful in the context of two-dimensional conformal field theory.
At times, the discussion will necessarily become technical in order to maintain precision. However, we will aim to motivate each step and provide clear explanations for the objects we introduce, helping the reader follow the structure of the construction, understand where we are headed, and see why each ingredient is introduced and how it fits into the overall picture.
Recall that in the previous section, we motivated the need to construct the cohomology of $\mathbb{R}\times\mathbb{C}\setminus\{0\}$ in order to obtain an object analogous to
(4.0.1) $H^{(0,0)}_{\bar{\partial}}\big((\mathbb{C}\setminus\{0\})\otimes\mathfrak{g}\big)={\mathbb{C}}[z,z^{-1}]\otimes\mathfrak{g}\,$,
which underlies the mode expansions of the Kac-Moody currents. In our case, however, we no longer work with the differential $\bar{\partial}$, but with a modified differential $\mathrm{d}^{\prime}=\bar{\partial}+\mathrm{d}_{\tau}$. Accordingly, the object we want to study is schematically of the form
(4.0.2) $H_{\mathrm{d}^{\prime}}\big((\mathbb{R}\times\mathbb{C}\setminus\{0\})\otimes\mathfrak{G}\big)\,$,
where $\mathfrak{G}=(\mathfrak{g}\xrightarrow{\mu_{1}}\mathfrak{h},\mu_{2})$ is the Lie 2-algebra used to define our three-dimensional theory (3.0.1). To construct such an object, we must first understand the structure of the differential complex defined by $\mathrm{d}^{\prime}$.
4.1. The mixed de Rham-Dolbeault complexes
Let $M$ denote a 3d manifold equipped with a THF $\mathcal{F}$. Locally, there are coordinate charts $(\tau,z,\bar{z})\in{\mathbb{R}}\times{\mathbb{C}}$ on $M$ which fixes a canonical basis $\partial_{\tau}$ spanning the topological directions $T_{(\tau,z,\bar{z})}\mathcal{F}$. The complexified quotient bundle
(4.1.1) $Q_{\mathbb{C}}=TM/T\mathcal{F}\otimes{\mathbb{C}}\cong Q^{(1,0)}\otimes Q^{(0,1)}\,$,
then determines a projection $TM\otimes{\mathbb{C}}\rightarrow Q^{(1,0)}$ whose kernel
$V_{\mathcal{F}}=\operatorname{ker}(TM\otimes{\mathbb{C}}\rightarrow Q^{(1,0)})\subset TM\otimes{\mathbb{C}}$
defines a subbundle spanned locally by the vector fields $\partial_{\tau},\partial_{\bar{z}}$. In other words, we have at each point $(\tau,z,\bar{z})\in M$ the tangent spaces
(4.1.2) $V_{\mathcal{F}}|_{(\tau,z,\bar{z})}=\mathrm{span}(\partial_{\tau},\partial_{\bar{z}})\,,\quad Q^{(1,0)}|_{(\tau,z,\bar{z})}=\mathrm{span}(\partial_{z})\,.$
The space $V_{\mathcal{F}}$ plays a crucial role in the characterization of manifolds with THFs $(M,\mathcal{F})$ [20]. Moreover, it also gives rise to a local decomposition of the complexified de Rham complex $\Omega_{\mathrm{dR}}^{\bullet}(M)$ on $M$, such that the exterior derivative splits
(4.1.3) ${\mathrm{d}}={\mathrm{d}}^{\prime}+\partial\quad\text{on}\quad\mathcal{A}^{p,q}=\wedge^{p}(V_{\mathcal{F}}^{*})\otimes\wedge^{q}(Q^{(1,0)})^{*},$
into the holomorphic Dolbeault differential $\partial$ plus the differential $\mathrm{d}$$^{\prime}$ on the so-called $\mathcal{F}$-basic holomorpic forms [^5] More precisely, these are sheaves of holomorphic functions $f$ on $M$ such that $L_{X}f=0$ for all $X\in V_{F}$.. In this decomposition $V_{\mathcal{F}}^{*}$ is locally spanned by ${\mathrm{d}}\bar{z},{\mathrm{d}}\tau$, while $\big(Q^{(1,0)}\big)^{*}$ is locally spanned by ${\mathrm{d}}z$. Note, for reference, that $\mathcal{A}^{p,q}$ is what we usually denote by $\Omega^{(p,q)}(M)$. Since the space $\mathcal{A}^{p,q}$ will play an important role, let us provide explicit examples of its elements for the lowest degrees.
The space $\mathcal{A}^{0,0}$ consists of smooth complex-valued functions. An element of $\mathcal{A}^{1,0}$ can be expressed locally as
(4.1.4) $L_{\bar{z}}(z,\tau)\,\mathrm{d}\bar{z}+L_{\tau}(z,\tau)\,\mathrm{d}\tau\,.$
At degree 2, elements of $\mathcal{A}^{1,1}$ are locally given by
(4.1.5) $\tilde{H}_{z\bar{z}}(z,\tau)\,\mathrm{d}z\wedge\mathrm{d}\bar{z}+\tilde{H}_{z\tau}(z,\tau)\,\mathrm{d}z\wedge\mathrm{d}\tau\,$,
while elements of $\mathcal{A}^{2,0}$ look like
(4.1.6) $H_{\tau\bar{z}}(z,\tau)\,\mathrm{d}\tau\wedge\mathrm{d}\bar{z}\,.$
In particular, the differential $\mathrm{d}$$^{\prime}$ raises the $p$-degree by 1, while $\partial$ raises the $q$-degree by 1.
Remark 1:
If $M$ is split, that is, if $M\cong\Sigma\times S$ with $\Sigma$ a Riemann surface and $S$ a $1$-manifold, then there is an isomorphism of cochain complexes
(4.1.7) $\left(\mathcal{A}^{\bullet,q},{\mathrm{d}}^{\prime}\right)\cong\left(\Omega^{(q,\bullet)}_{\mathrm{Dol}}(\Sigma)\otimes\Omega^{\bullet}_{\mathrm{dR}}(S),\bar{\partial}\otimes\mathbb{1}+\mathbb{1}\otimes{\mathrm{d}}_{\tau}\right)$
for $q=0,1$, where $\Omega^{(q,\bullet)}_{\mathrm{Dol}}(\Sigma)$ is the Dolbeault complex and $\Omega^{\bullet}_{\mathrm{dR}}(S)$ the de Rham complex. In this case, we can also define a different complex
(4.1.8) $(\mathcal{B}^{\bullet,q},\partial+\bar{\partial})=\left(\Omega^{(\bullet,\bullet)}_{\mathrm{Dol}}(\Sigma)\otimes\Omega^{q}_{\mathrm{dR}}(S),(\partial+\bar{\partial})\otimes\mathbb{1}\right)$
for $q=0,1$. In other words, the idea is that the differential in $\mathcal{A}$ can be thought of as split into $\mathrm{d}=\mathrm{d}^{\prime}+\partial$, while in $\mathcal{B}$ it splits as $\mathrm{d}=(\partial+\bar{\partial})+\mathrm{d}_{\tau}$. Crucially, the complexes $\mathcal{A}^{\bullet,q}$ and $\mathcal{B}^{\bullet,q}$ are not quasi-isomorphic, but there do exist isomorphisms
(4.1.9) $\left(\bigoplus_{q=0,1}\left(\mathcal{A}^{\bullet,q},{\mathrm{d}}^{\prime}\right),\partial\right)\cong\Omega_{\mathbb{C}}^{\bullet}(M)\cong\left(\bigoplus_{q=0,1}\left(\mathcal{B}^{\bullet,q},(\partial+\bar{\partial})\right),{\mathrm{d}}_{\tau}\right)\,$,
when we add back in the ”missing” differentials in $\mathcal{A},\mathcal{B}$. This gives a precise notion of the chirality vector $\ell$ introduced in [19]: it specifies which differential is “missing”.
4.1.1. Tensoring with $\mathfrak{G}$
At the beginning of this section, we stated that our goal was to compute the cohomology of $M\otimes\mathfrak{G}$. However, unlike the 2d case, $\mathfrak{G}=(\mathfrak{g}\xrightarrow{\mu_{1}}\mathfrak{h},\mu_{2})$ itself is a chain complex with differential $\mu_{1}$. Consequently, we are actually dealing with a tensor product complex of the form
(4.1.10) $\cdots$$\mathcal{A}^{0,q}\otimes\mathfrak{h}$$\mathcal{A}^{1,q}\otimes\mathfrak{h}$$\cdots$$\cdots$$\mathcal{A}^{0,q}\otimes\mathfrak{g}$$\mathcal{A}^{1,q}\otimes\mathfrak{g}$$\cdots$$\scriptstyle{{\mathrm{d}}^{\prime}}$$\scriptstyle{\mu_{1}}$$\scriptstyle{\mu_{1}}$$\scriptstyle{{\mathrm{d}}^{\prime}}$
for $q=0,1$. As a tensor product complex $\mathcal{A}^{\bullet,q}\otimes\mathfrak{G}$ is equipped with the differential
(4.1.11) $\hat{\mathrm{d}}^{\prime}={\mathrm{d}}^{\prime}\otimes\mathbb{1}-(-1)^{\text{deg}}\mathbb{1}\otimes\mu_{1}\,$,
where “deg” denotes the homogeneous degree of the first entry. More explicitly, for each $\alpha\otimes\mathsf{x}\in\mathcal{A}^{\bullet,q}\otimes\mathfrak{G}$ we have
(4.1.12) $\hat{\mathrm{d}}^{\prime}(\alpha\otimes\mathsf{x})={\mathrm{d}}^{\prime}\alpha\otimes\mathsf{x}-(-1)^{|\alpha|}\alpha\otimes\mu_{1}(\mathsf{x})\,$,
where $|\alpha|$ is the homogeneous degree of $\alpha\in\mathcal{A}^{\bullet,q}$.
The crucial observation is then that, as we shall now discuss, the currents associated to the symmetry of the three-dimensional theory determine precisely the cohomology classes in the tensor product complex $\mathcal{A}^{\bullet,q}\otimes\mathfrak{G}$!
4.2. Revisiting the Three-Dimensional Theory
Let us now apply the framework developed above to reformulate our 3d topological-holomorphic field theory $S[g,\Theta]$ in terms of the tensor product complex $\mathcal{A}^{p,q}\otimes\mathfrak{G}$.
We begin by noting that, since the component $\Theta_{z}$ is missing from the action (3.0.1), we can view the fundamental fields of $S[g,\Theta]$ as elements
$g\in\mathcal{A}^{0,0}\otimes G,\quad\Theta\in\mathcal{A}^{1,0}\otimes\mathfrak{h}\,.$
This then allows us to write $S[g,\Theta]$ in a very compact form
(4.2.1) $S[g,\Theta]=\int_{M}\langle\partial({\mathrm{d}}^{\prime}gg^{-1}),g\rhd\Theta\rangle-\tfrac{1}{2}\langle\mu_{1}(\partial(g\rhd\Theta)),g\rhd\Theta\rangle\,.$
The symmetry transformation parameters of (3.0.2) and (3.0.4) can also be identified with elements of the complex as
(4.2.2) $\alpha,\tilde{\alpha}\in\mathcal{A}^{0,0}\otimes\mathfrak{g}\,,\quad\Gamma,\tilde{\Gamma}\in\mathcal{A}^{1,0}\otimes\mathfrak{h}\,$,
as well as the currents
(4.2.3) $\tilde{L}=g^{-1}\partial g\,,\quad\tilde{H}=\partial\Theta+\mu_{2}(g^{-1}\partial g,\Theta)\,$,
and
(4.2.4) $L=-\mathrm{d}^{\prime}g\,g^{-1}-\mu_{1}(\Theta)\,,\quad H=g\rhd\left(\mathrm{d}^{\prime}\Theta-\tfrac{1}{2}[\Theta,\Theta]\right)\,$,
which can be written as
(4.2.5) $L\in\mathcal{A}^{1,0}\otimes\mathfrak{g}\,,\quad H\in\mathcal{A}^{2,0}\otimes\mathfrak{h}$
(4.2.6) $\tilde{L}\in\mathcal{A}^{0,1}\otimes\mathfrak{g}\,,\quad\tilde{H}\in\mathcal{A}^{1,1}\otimes\mathfrak{h}\,.$
We are now in a position to formulate a pair of conservation conditions for our three-dimensional theory in a form analogous to the 2d WZW case $\bar{\partial}J=0$ and $\partial\bar{J}=0$.
[…]
(4.2.7) $\hat{\mathrm{d}}^{\prime}(\tilde{L},\tilde{H})=({\mathrm{d}}^{\prime}\tilde{L}-\mu_{1}(\tilde{H}),{\mathrm{d}}^{\prime}\tilde{H})=0\,$,
[…]
(4.2.8) $\hat{\mathrm{d}}^{\prime}(\tilde{\alpha},\tilde{\Gamma})=({\mathrm{d}}^{\prime}\tilde{\alpha}+\mu_{1}(\tilde{\Gamma}),{\mathrm{d}}^{\prime}\tilde{\Gamma})=0\,$,
[…]
Conversely, the conservation equations for the currents $(L,H)$ given in (3.0.2) associated to the left symmetry of (4.2.1) are given by
(4.2.9) $\partial(L,H)=(\partial L,\partial H)=0\,$,
so that $(L,H)$ is a representative of a cohomology class in $H^{(\bullet,1)}_{\partial}(M\otimes\mathfrak{G})$, whereas the differential constraints (3.0.5) take the form
(4.2.10) $\partial(\alpha,\Gamma)=(\partial\alpha,\partial\Gamma)=0\,$,
so that $(\alpha,\Gamma)$ is a representative of a cohomology class in $H^{(\bullet,0)}_{\partial}(M\otimes\mathfrak{G})$.
[…]
It is important to note that in this case, the chiral and anti-chiral sectors are not “symmetric” to one another in an obvious way. In particular, as we will see, only the chiral sector will give rise to the infinite dimensional symmetry algebra. Consequently, we will focus on the cohomology group
(4.2.11) $H_{\hat{\mathrm{d}}^{\prime}}^{(\bullet,q)}(M\otimes\mathfrak{G})\,$,
[…]
Remark 2:
In [19, 40], a non-chiral version of the three-dimensional field theory was also considered. This variant is formulated using the decomposition of the de Rham differential associated with the bicomplex $\mathcal{B}$$^{\bullet,\bullet}$ defined in (4.1.8), given by the splitting
(4.2.12) $\mathrm{d}=(\partial+\bar{\partial})+\mathrm{d}_{\tau}\,.$
The fields consist of a group-valued field $g\in{\mathcal{B}}^{0,0}\otimes G$ and a one-form field $\Theta\in{\mathcal{B}}^{1,0}\otimes\mathfrak{h}$. The corresponding action is given by
(4.2.13) $S_{\text{nc}}[g,\Theta]=\int_{M}\left\langle\mathrm{d}_{\tau}(\mathrm{d}_{\Sigma}gg^{-1}),g\rhd\Theta\right\rangle-\tfrac{1}{2}\left\langle\mu_{1}\big(\mathrm{d}_{\tau}(g\rhd\Theta)\big),g\rhd\Theta\right\rangle\,$,
where we have written ${\mathrm{d}}_{\Sigma}=\partial+\bar{\partial}$ for short. This theory also features left- and right-acting symmetries, with associated conserved currents $(L_{\text{nc}},H_{\text{nc}})$ and $(\tilde{L}_{\text{nc}},\tilde{H}_{\text{nc}})$, which satisfy explicitly non-chiral conservation equations (see equation (6.3) in [40]):
${\mathrm{d}}_{\tau}(L_{\text{nc}},H_{\text{nc}})=0\,,\qquad({\mathrm{d}}_{\Sigma}-\mu_{1})(\tilde{L}_{\text{nc}},\tilde{H}_{\text{nc}})=0\,.$
This highlights a fundamental difference between the non-chiral theory $S_{\mathrm{nc}}$ and its chiral counterpart: their dynamics are fundamentally different — indeed, $S_{\mathrm{nc}}$ is fully topological (see Theorem 6.2 in [19]).
4.3. Mode expansion of the currents
[…]
(4.3.1) $H_{\hat{\mathrm{d}}^{\prime}}^{(\bullet,q)}(M\otimes\mathfrak{G})\,$,
with $q=0,1$. Our next goal is to obtain an explicit expression for this cohomology group. We will proceed in two steps: first, we will describe $H^{(\bullet,q)}_{\mathrm{d}^{\prime}}(M)$ following [22], and then we will apply the Künneth formula to compute the cohomology of the tensor product complex.
Similarly than in the two-dimensional pure Dolbeault case, there is a linear isomorphism
(4.3.2) ${\mathrm{d}}z\wedge-:H^{(\bullet,0)}_{\mathrm{d}^{\prime}}(M)\to H^{(\bullet,1)}_{\mathrm{d}^{\prime}}(M)\,.$
Hence, it suffices to find an explicit realization of $H^{(\bullet,0)}_{\mathrm{d}^{\prime}}(M)$, since elements in $H^{(\bullet,1)}_{\mathrm{d}^{\prime}}(M)$ can be obtained by wedging elements of the former with $\mathrm{d}z$. The cohomology of $(\mathcal{A}^{\bullet,0},\mathrm{d}^{\prime})$ was completely characterized in [22] and is concentrated in degrees zero and one. The zeroth cohomology consists of polynomials in the holomorphic variable $z$, that is,
(4.3.3) $H^{(0,0)}_{\mathrm{d}^{\prime}}(M)=\mathbb{C}[z]\,.$
Unlike in the Dolbeault complex, no negative powers of $z$ appear here. Consequently, any element $\tilde{\alpha}\in H^{(0,0)}_{\mathrm{d}^{\prime}}(M)$ can be expanded in modes as
(4.3.4) $\tilde{\alpha}=\sum_{n=0}^{\infty}\tilde{\alpha}_{n}z^{n}\,.$
The degree-one cohomology $H^{(1,0)}_{\mathrm{d}^{\prime}}(M)$ is more subtle and requires introducing additional structures. Since the negative powers of $z$ that appear in the holomorphic setting are absent from $H^{(0,0)}_{\mathrm{d}^{\prime}}(M)$, the idea is that in the raviolo framework, the role of $z^{-1}$ is played by a different object with analogous properties, which appears in $H^{(1,0)}_{\mathrm{d}^{\prime}}(M)$. Concretely, given a point $\underline{z}=(\tau,z)\in{\mathbb{R}}\times{\mathbb{C}}$, this is given by the $(1,0)$-form
(4.3.5) $\omega(\underline{z})=\frac{\tau\,{\mathrm{d}}\bar{z}-2\bar{z}\,{\mathrm{d}}\tau}{(|z|^{2}+\tau^{2})^{3}}\in V^{\ast}_{{\mathcal{F}}}\,.$
[…]
The key property of $\omega$ is an analogue of Cauchy’s residue theorem: its integral over a two-sphere centered at the origin is given by
(4.3.6) $\oint_{S^{2}}{\mathrm{d}}z\wedge\omega=8\pi i\,.$
Moreover, just as $z^{-1}$ is the Green’s function for $\partial$ on $\mathbb{C}\setminus\{0\}$, the $(1,0)$-form $\omega$ serves as the Green’s function for $\mathrm{d}$$^{\prime}$ on $\mathbb{R}\times\mathbb{C}\setminus\{0\}$. Consequently, they both act as propagators in field theories with the corresponding kinetic operators.
We can now use $\omega$ to give an explicit description of $H^{(1,0)}_{\mathrm{d}^{\prime}}(M)$. First, we introduce the so-called degree-one raviolo differential forms, which are obtained by taking derivatives of $\omega$ with respect to $z$. Specifically, we define
(4.3.7) $\Omega^{m}=\frac{(-1)^{m}}{m!}\partial_{z}^{m}\omega\,.$
Each $\Omega^{m}$ is a one-form, and they obey the relations
(4.3.8) $z^{n}\Omega^{m}=\begin{cases}0&\text{for }n>m\\ \Omega^{m-n}&\text{for }n\leq m\end{cases}\,$,
together with the holomorphic derivative identity [22]
(4.3.9) $\partial_{z}\Omega^{m}=-(m+1)\Omega^{m+1}\,.$
In other words, the raviolo differential forms serve a role analogous to that of negative modes in ${\mathbb{R}}\times{\mathbb{C}}\setminus\{0\}$, but in the form of 1-forms with non-trivial degree. This is consistent with the expectation from Hartog’s theorem.
The degree-one cohomology is then given by [22]
(4.3.10) $H^{(1,0)}_{\mathrm{d}^{\prime}}(M)=\mathrm{span}_{\mathbb{C}}\{\Omega^{0},\Omega^{1},\Omega^{2},\dots\}\,$,
so that any element $\tilde{\Gamma}\in H^{(1,0)}_{\mathrm{d}^{\prime}}(M)$ can be expanded in modes as
(4.3.11) $\tilde{\Gamma}=\sum_{m=0}^{\infty}\tilde{\Gamma}_{m}\Omega^{m}\,.$
We will denote the full cohomology by
(4.3.12) $\mathcal{K}^{\bullet}_{\text{poly}}:=H^{(\bullet,0)}_{\mathrm{d}^{\prime}}(M)\,$,
and refer to it as the space of polynomials on the raviolo. This space is the analogue of the Laurent polynomials $\mathbb{C}[z,z^{-1}]$ on $\mathbb{C}^{\times}$ and provides the appropriate algebraic structure for constructing the mode expansions, ultimately leading to the centrally extended affine graded Lie algebra.
Having found an explicit realization for $H^{(\bullet,0)}_{\mathrm{d}^{\prime}}(M)$, we can use the linear isomorphism in (4.3.2) to construct elements in $H^{(\bullet,1)}_{\mathrm{d}^{\prime}}(M)$. The existence of this isomorphism implies that any element $\tilde{L}\in H^{(0,1)}_{\mathrm{d}^{\prime}}(M)$ can be written as $\tilde{L}=\tilde{L}_{z}\mathrm{d}z$ for some $\tilde{L}_{z}\in H^{(0,0)}_{\mathrm{d}^{\prime}}(M)$ and similarly with elements in $H^{(1,1)}_{\mathrm{d}^{\prime}}(M)$. With a slight abuse of notation we will write
(4.3.13) $\mathrm{d}z\wedge\mathcal{K}_{\mathrm{poly}}^{\bullet}:=H^{(\bullet,1)}_{\mathrm{d}^{\prime}}(M)\,.$
With an explicit expression for both $H^{(\bullet,0)}_{\mathrm{d}^{\prime}}(M)$ and $H^{(\bullet,1)}_{\mathrm{d}^{\prime}}(M)$ at hand, we can proceed to determine the cohomology of the tensor product complex $H_{\hat{\mathrm{d}}^{\prime}}^{(\bullet,q)}(M\otimes\mathfrak{G})$. To do so, we make use of the Künneth formula which relates the cohomology of the tensor product with the tensor product of the cohomologies
(4.3.14) $H_{\hat{\mathrm{d}}^{\prime}}^{(0,q)}(M\otimes\mathfrak{G})\cong\bigoplus_{0=m+k}H^{(m,q)}_{{\mathrm{d}}^{\prime}}(M)\otimes H^{k}_{\mu_{1}}(\mathfrak{G})\,$,
where the ”correction” to the Künneth formula, which are the derived Ext-groups, do not appear here as both of our complexes ${\mathcal{A}},\mathfrak{G}$ are free.
Since $\mathfrak{G}$ is concentrated in degrees 1-1 and 0, its cohomology will also be concentrated in these degrees. Let $H^{\bullet}(\mathfrak{G})=V\oplus\mathfrak{n}$ denote the cohomology of the Lie 2-algebra $\mathfrak{G}$, with
(4.3.15) $V=\ker\mu_{1}=H^{-1}(\mathfrak{G}),\qquad\mathfrak{n}=\mathfrak{g}/\mathrm{im}(\mu_{1})=H^{0}(\mathfrak{G})\,$,
where $V$ is an Abelian $\mathfrak{n}$-module. By a theorem of Gerstenhaber, the cohomology $H^{\bullet}(\mathfrak{G})$ is part of the data which characterizes Lie 2-algebras $\mathfrak{G}$ up to equivalence [41]. In the special case where $\mu_{1}=0$, namely when $\mathfrak{G}$ is called skeletal, then $\mathfrak{G}=H^{\bullet}(\mathfrak{G})$ is its own cohomology. On the other hand, if $\mu_{1}=\mathbb{1}$ is the identity, then $H^{\bullet}(\mathfrak{G})=0$ is trivial.
We then have the following theorem.
4.3.1 Theorem.
[…]
(4.3.16) $(\tilde{\alpha},\tilde{\Gamma})\in H^{(0,0)}_{\hat{\mathrm{d}}^{\prime}}(M\otimes\mathfrak{G})=(\mathcal{K}_{\mathrm{poly}}^{0}\otimes\mathfrak{n})\oplus(\mathcal{K}^{1}_{\mathrm{poly}}\otimes V)\,.$
[…]
(4.3.17) $(\tilde{L},\tilde{H})\in H_{\hat{\mathrm{d}}^{\prime}}^{(0,1)}(M\otimes\mathfrak{G})=(({\mathrm{d}}z\wedge\mathcal{K}_{\mathrm{poly}}^{0})\otimes\mathfrak{n})\oplus(({\mathrm{d}}z\wedge\mathcal{K}_{\mathrm{poly}}^{1})\otimes V)$
Proof.
The proof follows immediately from equations (4.2.7) and (4.2.8) ∎
The fundamental consequence of this theorem is that it provides a mode expansion for both our symmetry transformation parameters and higher currents. Specifically, we can express them as
(4.3.18) $\tilde{\alpha}=\sum_{n=0}^{\infty}\tilde{\alpha}_{n}z^{n}\,,\quad\tilde{\Gamma}=\sum_{m=0}^{\infty}\tilde{\Gamma}_{m}\Omega^{m}$
with $\tilde{\alpha}_{n}\in\mathfrak{n}$ and $\tilde{\Gamma}_{m}\in V$. Similarly, the higher currents admit the expansions
(4.3.19) $\tilde{L}=\sum_{n=0}^{\infty}\tilde{L}_{n}z^{n}\,{\mathrm{d}}z\,,\quad\tilde{H}={\mathrm{d}}z\wedge\sum_{m=0}^{\infty}\tilde{H}_{m}\Omega^{m}$
where $\tilde{L}_{n}\in\mathfrak{n}$ are $\mathfrak{n}$ and $\tilde{H}_{m}\in V$. In the rest of the paper, we will work exclusively in cohomology. In this setting the fields depend only on the holomorphic coordinate $z$. For functions, this is immediate because every cohomology class is represented by a Taylor series in $z$. For $(1,0)$-forms, recall that $\partial_{\tau}\omega$ and $\partial_{\bar{z}}\omega$ are $\mathrm{d}^{\prime}$-exact, which implies that $\partial_{\tau}\Omega^{m}$ and $\partial_{\bar{z}}\Omega^{m}$ are also $\mathrm{d}^{\prime}$-exact for every $m$. Therefore, the only nontrivial information in cohomology comes from the dependence on $z$.
5. Radial Quantisation
[…]
Before turning to the details, let us indicate why a very similar construction extends so smoothly to $\mathbb{R}$$\times{\mathbb{C}}$. In the complex plane two ingredients were essential: the mode expansion and Cauchy’s residue theorem, which allowed us to translate the singular behaviour of operator products into commutators. In the three-dimensional setting we have already established the mode expansion of the currents in (4.3.19), and we also introduced the $(1,0)$-form $\omega$ in (4.3.5), which satisfies
(5.0.1) $\oint_{S^{2}}\mathrm{d}z\wedge\omega=8\pi i\,.$
This provides a higher–dimensional analogue of Cauchy’s residue formula, allowing us to generalise statements involving circles $S^{1}\subset{\mathbb{C}}$ to spheres $S^{2}\subset{\mathbb{R}}\times{\mathbb{C}}$. In Euclidean signature one can again take the radial coordinate as time and, in direct analogy with the two–dimensional case, define a radial quantisation scheme. The generalised Cauchy formula then makes it possible to translate singular OPEs of the three–dimensional currents into commutation relations.
5.1. The Ward Identity
[…]
(5.1.1) ${\tilde{L}}=g^{-1}\partial_{z}g\,{\mathrm{d}}z\,,\qquad{\tilde{H}}=\mathrm{d}z\wedge\big(\partial_{z}\Theta+\mu_{2}(g^{-1}\partial_{z}g,\Theta)\big)\,$,
[…]
The goal is to compute the Ward Identity associated with the right acting transformation
(5.1.2) $g\to g+g{\tilde{\alpha}}\,,\quad\Theta\to\Theta-{\tilde{\alpha}}\rhd\Theta+{\tilde{\Gamma}}\,$.
Hence, we consider $\tilde{\alpha}$} and $\tilde{\Gamma}$} with compact support in a ball $B\subset{\mathbb{R}}\times{\mathbb{C}}$, vanishing outside $B$, and subject to the differential constraints
(5.1.3) $\mathrm{d}^{\prime}{\tilde{\alpha}}+\mu_{1}{\tilde{\Gamma}}=0\,,\quad\mathrm{d}^{\prime}{\tilde{\Gamma}}=0\,.$
Under this transformation, the action (4.2.1) varies as
[…]
where we have used (5.1.3) together with Stokes’ theorem.
Now in order to achieve a mode expansion as given in the previous section, we shall consider the fields and symmetry parameters appearing in the above equation as representatives of cohomology classes valued in $H^{\bullet}(\mathfrak{G})=V\oplus\mathfrak{n}$, [^7] This step is called a homotopy transfer (see appendix B). We shall come back to explain its physical meaning fully in a followup work. where, recall,
(5.1.5) $V=\ker(\mu_{1})\subset\mathfrak{h}\,,\qquad\mathfrak{n}=\mathfrak{g}/\operatorname{im}(\mu_{1})\,.$
To write expressions like $\langle\tilde{\alpha},\tilde{H}\rangle$ or $\tilde{\alpha},\tilde{\alpha}^{\prime}$], it is necessary to verify that these operations descend to cohomology. We include a proof of this fact in appendix B. For notational simplicity, we will use the same symbols [,][-,-], $\mu_{2}$, and $\langle-,$-$\rangle$ for the induced bracket, action, and bilinear form on $H^{\bullet}(\mathfrak{G})$. Moreover, given a basis $\{t_{a}\}$ of $\mathfrak{n}$ and $\{s_{b}\}$ of $V$, we shall use the same notation for the corresponding structure constants and bilinear form, whenever no confusion can arise:
(5.1.6) $t_{a},t_{b}]=f_{ab}^{c}\,t_{c}\,,\qquad\mu_{2}(t_{a},s_{b})=(\mu_{2})_{ab}^{c}\,s_{c}\,,\qquad\langle t_{a},s_{b}\rangle=\kappa_{ab}\,.$
This allows us to express the currents and transformation parameters as linear combinations of the generators of the algebras,
(5.1.7) ${\tilde{\alpha}}=\sum_{a=1}^{\dim\mathfrak{n}}{\tilde{\alpha}}^{a}t_{a}\,,\quad{\tilde{H}}=\sum_{b=1}^{\dim V}{\tilde{H}}^{b}s_{b}\,,\quad{\tilde{L}}=\sum_{a=1}^{\dim\mathfrak{n}}{\tilde{L}}^{a}t_{a}\,,\quad{\tilde{\Gamma}}=\sum_{b=1}^{\dim V}{\tilde{\Gamma}}^{b}s_{b}\,$,
and the Ward identity (2.1.1) takes the form [^8] Note that the symbol $\langle$-$\rangle$ denotes a correlation function, while the symbol $\langle-,$-$\rangle$ refers to the bilinear form on $\mathfrak{G}$.
(5.1.8) $\delta_{({\tilde{\alpha}},{\tilde{\Gamma}})}\langle X\rangle=\oint_{\partial B}{\mathrm{d}}z\wedge\left({\tilde{\alpha}}^{a}\langle{\tilde{H}}^{b}X\rangle+\langle{\tilde{L}}^{a}X\rangle{\tilde{\Gamma}}_{b}\right)\kappa_{ab}\,.$
We can now use this relation to determine the operator product expansion between the currents. To do so, we will make use the following lemma
5.1.1 Lemma.
Let and $S_{w}^{2}$ denote a ball centered at $w$. Then
(5.1.9) $\frac{1}{8\pi i}\oint_{S_{w}^{2}}\mathrm{d}z\wedge(z-w)^{n}\Omega_{z-w}^{m}=\delta_{m,n}$
where $\Omega_{z-w}^{m}=\frac{(-1)^{m}}{m!}\partial_{z}^{m}\omega(z-w)$ for $\omega$ defined in (4.3.5).
Proof.
The proof is an immediate consequence of the relation (4.3.8), together with the integral of $\omega$ over a sphere (5.0.1). ∎
We thus have the following theorem
5.1.2 Theorem.
The currents $\tilde{L}$} and $\tilde{H}$} have the following operator product expansions
(5.1.10) ${\tilde{L}}^{a}(z){\tilde{L}}^{b}(w)=\mathrm{reg}$
(5.1.11) ${\tilde{H}}^{a}(z){\tilde{L}}^{b}(w)=\Omega_{z-w}^{0}(\tilde{\mu}_{2})^{ab}_{c}{\tilde{L}}^{c}(w)+\Omega_{z-w}^{1}(\kappa^{-1})^{ab}+\mathrm{reg}$
(5.1.12) ${\tilde{H}}^{a}(z){\tilde{H}}^{b}(w)=\Omega_{z-w}^{0}\tilde{f}^{ab}_{c}{\tilde{H}}^{c}(w)+\mathrm{reg}$
where $\mathrm{reg}$ stand for terms which are regular as $z\to w$, and where we have defined the dual structure constants
(5.1.13) $\tilde{f}^{ab}_{c}=(\kappa^{-1})^{ax}(\kappa^{-1})^{by}f_{xy}^{z}\kappa_{zc},\qquad(\tilde{\mu}_{2})^{ab}_{c}=(\kappa^{-1})^{ax}\kappa_{cz}(\mu_{2})_{xy}^{z}(\kappa^{-1})^{yb}\,.$
Remark 3:
Note that the bilinear form $\langle\cdot,\cdot\rangle:\mathfrak{n}\otimes V\oplus V\otimes\mathfrak{n}\to\mathbb{C}$ pairs elements of $\mathfrak{n}$ with those of $V$, and thus, unlike the standard case, it does not provide an identification between either of these spaces and their respective duals. As a result, we cannot use $\kappa$ to raise or lower indices; this is why we define the dual structure constants $\tilde{f}^{ab}_{c}$ and $(\tilde{\mu}_{2})_{c}^{ab}$.
Proof.
We begin by computing the variations of the currents $\tilde{L}$} and $\tilde{H}$} under the symmetry transformation (5.1.2), which are given by
(5.1.14) $\delta_{(\tilde{\alpha},\tilde{\Gamma})}\tilde{L}$ $=[\tilde{L},{\tilde{\alpha}}]+\partial_{z}\tilde{\alpha}\,$,
(5.1.15) $\delta_{(\tilde{\alpha},\tilde{\Gamma})}\tilde{H}$ $=-\mu_{2}(\tilde{\alpha},\tilde{H})+\mu_{2}(\tilde{L},\tilde{\Gamma})+\partial\tilde{\Gamma}\,$.
taking the product of fields to be $X={\tilde{L}}^{b}(w)X^{\prime}$, with $X^{\prime}=\prod_{i}\mathcal{O}_{i}(u_{i})$ where $w\in B$ and all $u_{i}\notin B$, where the $\mathcal{O}_{i}$ denote arbitrary operator insertions, the vanishing of $\tilde{\alpha}$} and $\tilde{\Gamma}$} outside $B$ implies that $\delta_{{\tilde{\alpha}},{\tilde{\Gamma}}}$ acts only on ${\tilde{L}}^{b}(w)$. The Ward identity takes the form
(5.1.16) $\langle([\tilde{L},{\tilde{\alpha}}]^{b}+\partial_{z}\tilde{\alpha}^{b})X^{\prime}\rangle=\frac{1}{8\pi i}\oint_{\partial B}\mathrm{d}z\wedge{\tilde{\alpha}}^{c}\,\langle{\tilde{H}}^{a}(z){\tilde{L}}^{b}(w)X^{\prime}\rangle\kappa_{ca}\\ +\frac{1}{8\pi i}\oint_{\partial B}\mathrm{d}z\wedge\langle{\tilde{L}}^{a}(z){\tilde{L}}^{b}(w)X^{\prime}\rangle{\tilde{\Gamma}}^{c}\kappa_{ca}\,.$
Since the left-hand side contains no terms proportional to $\tilde{\Gamma}$}, we conclude
(5.1.17) ${\tilde{L}}^{a}(z){\tilde{L}}^{b}(w)=\mathrm{reg}\,.$
Let us now check that the OPE (5.1.11) reproduces the remaining terms. Inserting the first contribution from (5.1.11) in (5.1.16) yields
(5.1.18) $\frac{1}{8\pi i}\oint_{\partial B}\mathrm{d}z\wedge{\tilde{\alpha}}^{c}\Omega_{z-w}^{0}(\tilde{\mu}_{2})^{ab}_{d}\langle{\tilde{L}}^{d}(w)X^{\prime}\rangle\kappa_{ca}={\tilde{\alpha}}^{c}(\tilde{\mu}_{2})^{ab}_{d}\langle{\tilde{L}}^{d}(w)X^{\prime}\rangle\kappa_{ca}\,$,
where we have used the generalized Cauchy formula (5.1.9). On the other hand, the first term on the left-hand side of (5.1.16) reads in components
(5.1.19) ${\tilde{\alpha}}^{c}f^{b}_{dc}\langle{\tilde{L}}^{d}(w)X^{\prime}\rangle\,.$
To compare the two expressions we invoke the ad-invariance of the bilinear form,
(5.1.20) $\langle[t_{a},t_{b}],s_{c}\rangle=\langle t_{a},\mu_{2}(t_{b},s_{c})\rangle\quad\Rightarrow\quad f_{ab}^{d}\kappa_{dc}=\kappa_{ad}(\mu_{2})_{bc}^{d}\,$,
which implies
(5.1.21) $(\tilde{\mu}_{2})^{ab}_{d}\kappa_{ca}=(\kappa^{-1})^{ax}\kappa_{dz}(\mu_{2})_{xy}^{z}(\kappa^{-1})^{yb}\kappa_{ca}=f^{z}_{dc}\kappa_{zy}(\kappa^{-1})^{yb}=f^{b}_{dc}\,$,
where in the first line equality used the definition of $(\tilde{\mu}_{2})^{ab}_{d}$ given in (5.1.13), and in the third, the $\mathrm{ad}$-invariance of the bilinear form. Hence the contributions (5.1.18) and (5.1.19) agree. Similarly, if we insert the second contribution from (5.1.11) in (5.1.16) we find
(5.1.22) $\begin{split}\frac{1}{8\pi i}\oint_{\partial B}\mathrm{d}z\wedge{\tilde{\alpha}}^{c}\Omega_{z-w}^{1}(\kappa^{-1})^{ab}\langle X^{\prime}\rangle\kappa_{ca}&=\frac{1}{8\pi i}\oint_{\partial B}\mathrm{d}z\wedge\partial_{z}{\tilde{\alpha}}^{b}\Omega_{z-w}^{0}\langle X^{\prime}\rangle\\ &=\partial_{z}{\tilde{\alpha}}^{b}\langle X^{\prime}\rangle\end{split}$
where we used the holomorphic derivative relation $\Omega_{z-w}^{1}=-\partial_{z}\Omega_{z-w}^{0}$ to integrate by parts in the first step, and then applied the generalized Cauchy formula (5.1.9) in the second. This reproduces exactly the second term on the left-hand side of (5.1.16).
With a completely analogous computation, one verifies that the ${\tilde{H}}{\tilde{H}}-$OPE takes the form (5.1.12).
5.2. Commutation Relations
As we mentioned at the beginning of this section, the essence of this construction is that in analogy to the two-dimensional case, we can also perform radial quantisation. We choose our time coordinate to be the radial coordinate on $\mathbb{R}\times{\mathbb{C}}\cong{\mathbb{R}}^{3}$ so that equal-time slices will be given by spheres of fixed radius. Correlation functionas are then identified as vacuum expectation values of radially ordered products of operators
(5.2.1) $\langle a(z)b(w)\rangle=\langle 0|\mathcal{R}(a(z)b(w))|0\rangle\,$,
where the radial-ordering operator $\mathcal{R}$ is defined by
(5.2.2) $\mathcal{R}(a(z)b(w))=\begin{cases}\begin{aligned} &a(z)b(w)\quad\text{if}\quad|z|>|w|\\ &b(w)a(z)\quad\text{if}\quad|w|>|z|\end{aligned}\end{cases}\,.$
With this identification, we can relate the OPE with commutators. To this end, we consider the integral over the sphere
(5.2.3) $\oint_{S_{w}^{2}}\mathrm{d}z\wedge\,\mathcal{R}(a(z)b(w))\,$,
Next, we split the integration contour into two concentric, fixed–radius spheres: an outer one $B_{1}$ with $|z|>|w|$ and an inner one $B_{2}$ with $|z|<|w|$ with the same orientation [^9] More precisely, up to homotopy, $b^{1}$ can be written as the connected summation of $S_{w}^{2}$ and $B_{2}$, and the contribution from the attaching cylinder can be shown to vanish.. Radial ordering then implies
(5.2.4) $\begin{split}\oint_{S_{w}^{2}}\mathrm{d}z\wedge\mathcal{R}(a(z)b(w))&=\oint_{B_{1}}\mathrm{d}z\,a(z)b(w)-\oint_{B_{2}}\mathrm{d}z\,b(w)a(z)\\ &=[A,b(w)]\end{split}$
where we have defined the operator
(5.2.5) $A=\oint\mathrm{d}z\wedge a(z)\,.$
If we want this relation to hold as an operator identity, we must allow for an arbitrary number of additional fields to appear alongside $a(z)$ and $b(w)$ in a general correlator. In particular, the decomposition of the contour into $b^{1}$ and $B_{2}$ is justified only when $b(w)$ is the unique insertion with a singular OPE with $a(z)$ lying between the two spheres.
By integrating equation (5.2.4), we obtain the commutator between two operators, each defined as the contour integral of a holomorphic field:
(5.2.6) $A,B]=\oint_{S_{0}^{2}}\mathrm{d}w\oint_{S_{w}^{2}}\mathrm{d}z\wedge\mathcal{R}(a(z)b(w))\,.$
From the above relation, we can compute the commutation relations satisfied by the modes of the currents, which are summarized in the following theorem:
5.2.1 Theorem.
The commutation relations for the modes ${\tilde{L}}_{n}^{a}$, ${\tilde{H}}_{m}^{a}$ of the currents corresponding to the topological-holomorphic symmetry of the three-dimensional action (4.2.1) are given by
(5.2.7) ${\tilde{H}}_{n}^{a},{\tilde{H}}_{m}^{b}]=\tilde{f}^{ab}_{c}{\tilde{H}}_{n+m}^{c}\,,\quad[{\tilde{L}}_{n},{\tilde{L}}_{m}]=0\,$,
(5.2.8) ${\tilde{H}}_{n}^{a},{\tilde{L}}_{m}^{b}]=\begin{cases}(\tilde{\mu}_{2})^{ab}_{c}L_{m-n}^{c}+n(\kappa^{-1})^{ab}\delta_{n-1,m}\quad\text{if}\quad m\geq n-1\\ 0\qquad\qquad\qquad\qquad\qquad\quad\qquad\text{otherwise}\end{cases}\,$,
with the dual structure constants $\tilde{f}^{ab}_{c}$ and $(\tilde{\mu}_{2})^{ab}_{c}$ defined in (5.1.13).
Proof.
We recall the mode expansions of the components of the currents in a basis of the Lie algebras $\mathfrak{n}$ and $V$ respectively
(5.2.9) ${\tilde{L}}^{a}(z)=\sum_{n=0}^{\infty}{\tilde{L}}^{a}_{n}z^{n}\,,\quad{\tilde{H}}^{a}(z)=\sum_{m=0}^{\infty}{\tilde{H}}^{a}_{m}\Omega^{m}\,$,
where the modes can be written in terms of the components of the currents as
(5.2.10) ${\tilde{L}}_{n}^{a}=\frac{1}{8\pi i}\oint_{S^{2}}{\mathrm{d}}z\wedge\Omega^{n}{\tilde{L}}^{a}(z)\,,\quad{\tilde{H}}^{a}_{n}=\frac{1}{8\pi i}\oint_{S^{2}}{\mathrm{d}}z\wedge z^{n}{\tilde{H}}^{a}(z)\,.$
Then, using the definition of the commutator (5.2.6) we can write
(5.2.11) $\begin{split}[{\tilde{H}}^{a}_{n},{\tilde{H}}^{b}_{m}]&=\frac{1}{(8\pi i)^{2}}\oint_{S_{0}^{2}}{\mathrm{d}}w\oint_{S_{w}^{2}}{\mathrm{d}}z\wedge\,w^{n}\,z^{m}\,\mathcal{R}({\tilde{H}}^{a}(z){\tilde{H}}^{b}(w))\\ &=\frac{1}{(8\pi i)^{2}}\oint_{S_{0}^{2}}{\mathrm{d}}w\oint_{S_{w}^{2}}{\mathrm{d}}z\wedge\,w^{n}\,z^{m}\left(\Omega_{z-w}^{0}\tilde{f}^{ab}_{d}{\tilde{H}}^{d}(w)+\mathrm{reg}\right)\end{split}$
where in the second line we replaced the radially ordered product by the OPE (5.1.12), as the singular part is invariant under radial ordering due to the antisymmetry of $\tilde{f}^{ab}_{d}$ and only this part contributes to the integral. Expanding $z^{m}$ around $w$ and using the generalised Cauchy formula (5.1.9) we find
(5.2.12) ${\tilde{H}}^{a}_{n},{\tilde{H}}^{b}_{m}]=\frac{1}{8\pi i}\oint_{S_{0}^{2}}{\mathrm{d}}w\wedge\,w^{n+m}\,\tilde{f}^{ab}_{d}{\tilde{H}}^{d}(w)=f^{ab}_{d}{\tilde{H}}_{n+m}\,.$
Next, we have that ${\tilde{L}}_{n}^{a},{\tilde{L}}_{m}^{b}]=0$ since the ${\tilde{L}}{\tilde{L}}-$OPE is regular (5.1.10). Finally, for the mixed commutator we have
[…]
where in the second line we used the ${\tilde{H}}{\tilde{L}}-$OPE given in (5.1.11) and in the third line we expanded $z^{n}$ around $w$. The first term contributes only for $j=0$, whereas the second term will be non vanishing for $j=1$. Thus, using the generalized cauchy formula (5.1.9) we obtain
[…]
where we used the relation $\Omega_{w}^{m}w^{n}=\Omega_{w}^{m-n}$ for $m\geq n$, and the generalized Cauchy formula (5.1.9) for the second term.
Theorem 5.2.1 gives us commutation relations of the modes of the conserved currents associated to the infinite dimensional symemtry of the three-dimensional action (4.2.1). This is in complete analogy to the Kac-Moody algebra satisfied by the modes of the conserved currents in the Wess-Zumino-Witten model. In the next section, we will show that the current modes can be identified with the generators of a centrally extended affine graded Lie algebra
Remark 4:
In §4.3 of  [22], the authors construct a raviolo current algebra by tensoring a Lie superalgebra with the cohomology of the raviolo, and endow it with a bilinear form of even degree. In that setting, the vanishing brackets arise between generators valued in the odd part $\mathfrak{g}_{1}$ of the Lie superalgebra. In contrast, our construction uses a Lie 2-algebra tensored with the same cohomology, but equipped with a bilinear form of degree 1-1. As a consequence, the vanishing brackets in our case occur between degree 0 elements (the $L_{n}$’s), while the non-trivial brackets involve degree 1-1 elements (the $H_{n}$’s). This shift in grading structure can be traced back to the differing degrees of the bilinear forms used in each construction.
Remark 5:
The Noether currents of the non-chiral 3d action (4.2.13), defined in terms of $\mathcal{B}$}, was studied in §6.3 of [40]. The main result there was that the resulting differential graded commutator algebra — analogues of (5.2.7), (5.2.8) — was found to be closely related to a higher derived notion of Lax integrability, similar to the situation with the Wess-Zumino-Witten model [42]. The authors expect that these methods could similarly be applied to analyze the Lax integrability of the chiral theory (4.2.1).
6. The Centrally Extended Affine Graded Lie Algebra
In the previous section, we computed the commutation relations of the currents associated with the infinite-dimensional symmetry of the three-dimensional action (4.2.1). In this section, we show that these commutation relations define a centrally extended affine graded Lie algebra.
Kac-Moody Algebra
We begin by briefly recalling the construction in two dimensions, where the commutators of the WZW model give rise to a centrally extended affine Lie algebra. This infinite-dimensional algebra, known as the Kac–Moody algebra at level $k$, is defined as
(6.0.1) $\hat{\mathfrak{g}}_{k}:=\mathfrak{g}\otimes{\mathbb{C}}[z,z^{-1}]\oplus k\mathbb{C}\,$,
with Lie bracket
(6.0.2) $X\otimes f(z),Y\otimes g(z)]=[X,Y]\otimes(fg)(z)+k\langle X,Y\rangle\operatorname{Res}_{z=0}(f\partial_{z}g)\,$,
for $X\otimes f(z),Y\otimes g(z)\in\hat{\mathfrak{g}}_{k}$. The second term defines a nontrivial $2$-cocycle,
(6.0.3) $\omega(X\otimes f(z),Y\otimes g(z)):=k\langle X,Y\rangle\operatorname{Res}_{z=0}(f\partial_{z}g)\,$,
which satisfies the 2-cocycle condition
(6.0.4) $\omega\big([X\otimes f,Y\otimes g],Z\otimes h\big)+\omega\big([Y\otimes g,Z\otimes h],X\otimes f\big)\\ +\omega\big([Z\otimes h,X\otimes f],Y\otimes g\big)=0\,$,
for arbitrary elements $X\otimes f,Y\otimes g,Z\otimes h\in\hat{\mathfrak{g}}_{k}$. This relation holds provided that the bilinear form $\langle\cdot,\cdot\rangle$ on $\mathfrak{g}$ is invariant. The cocycle ensures that the central extension is compatible with the Lie algebra structure, guaranteeing that $\hat{\mathfrak{g}}_{k}$ satisfies the Jacobi identity and is thus a well-defined Lie algebra.
The identification between this realization of the Kac–Moody algebra and the algebra generated by the modes of the symmetry currents in the WZW model is given by
(6.0.5) $J_{n}^{a}\mapsto t_{a}\otimes z^{n}\in\mathfrak{g}\otimes\mathbb{C}[z,z^{-1}]\,$,
where $\{t_{a}\}$ is a basis of $\mathfrak{g}$. A direct computation then shows that the commutators of the currents in the WZW model
(6.0.6) $J_{n}^{a},J_{m}^{b}]=f_{c}^{ab}J_{n+m}^{c}+k\,\kappa^{ab}\delta_{n+m,0}$
match the Lie bracket (6.0.2) under the identification (6.0.5).
6.1. The Affine Graded Lie Algebra
The affine graded Lie algebra is constructed in close analogy with its lower-dimensional counterpart, ensuring that its commutation relations reproduce those of our three-dimensional theory.
To begin, we consider the dual vector spaces $\mathfrak{n}^{\vee}$ and $V^{\vee}$, equipped with dual bases $\{s^{a}\}$ and $\{t^{b}\}$, respectively:
(6.1.1) $\mathfrak{n}^{\vee}=\mathrm{span}(s^{a})\,,\qquad V^{\vee}=\mathrm{span}(t^{b})\,.$
These dual bases are constructed from the original bases $\{t_{a}\}$ of $\mathfrak{n}$ and $\{s_{b}\}$ of $V$, using the non-degenerate (invertible) bilinear form $\langle\cdot,\cdot\rangle$, as follows:
(6.1.2) $t^{a}=(\kappa^{-1})^{ab}\langle-,s_{b}\rangle\,,\qquad s^{a}=(\kappa^{-1})^{ab}\langle t_{a},-\rangle\,.$
By construction, these satisfy the duality conditions:
(6.1.3) $t^{a}(t_{b})=\delta^{a}_{b}\,,\qquad s^{a}(s_{b})=\delta^{a}_{b}\,.$
We endow the graded vector space $\mathfrak{G}^{\vee}=\mathfrak{n}^{\vee}\oplus V^{\vee}$ with the following brackets:
(6.1.4) $t^{a},t^{b}]^{\vee}=\tilde{f}^{ab}_{c}\,t^{c}\in V^{\vee},\qquad\mu_{2}^{\vee}(t^{a},s^{b})=(\tilde{\mu}_{2})^{ab}_{c}\,s^{c}\in\mathfrak{n}^{\vee}\,$,
where the structure constants are defined by
(6.1.5) $\tilde{f}^{ab}_{c}=(\kappa^{-1})^{ax}(\kappa^{-1})^{by}f_{xy}^{z}\kappa_{zc},\qquad(\tilde{\mu}_{2})^{ab}_{c}=(\kappa^{-1})^{ax}\kappa_{cz}(\mu_{2})_{xy}^{z}(\kappa^{-1})^{yb}\,$,
as in equation (5.1.13). We assign degree 0 to $V^{\vee}$ and degree 1-1 to $\mathfrak{n}^{\vee}$. With these conventions, we have the following:
6.1.1 Proposition.
The brackets defined in (6.1.4) satisfy the graded Jacobi identity. Hence, they endow $\mathfrak{G}^{\vee}$ with the structure of a graded Lie algebra.
Proof.
We consider
[…]
where we used the definition of the dual structure constants given in (6.1.5). Summing the three terms that appear in the graded Jacobi identity, we obtain
(6.1.7) $t^{a},[t^{b},t^{c}]^{\vee}]^{\vee}+[t^{c},[t^{a},t^{b}]^{\vee}]^{\vee}+[t^{b},[t^{c},t^{a}]^{\vee}]^{\vee}=\\ (\kappa^{-1})^{ar}(\kappa^{-1})^{bs}(\kappa^{-1})^{ct}\big(f_{ry}^{z}f_{st}^{y}+f_{ty}^{z}f_{rs}^{y}+f_{sy}^{z}f_{tr}^{y}\big)\kappa_{ze}t^{e}=0$
The expression in parentheses vanishes by the Jacobi identity for the Lie bracket of $\mathfrak{n}$, whose structure constants are given by $f_{ab}^{c}$. Next, for the mixed brackets involving both $t^{a}$ and $s^{b}$, we compute:
(6.1.8) $\begin{split}\mu_{2}^{\vee}(t^{a},\mu_{2}^{\vee}(t^{b},s^{c}))&=(\tilde{\mu}_{2})^{ad}_{e}(\tilde{\mu}_{2})^{bc}_{d}s^{e}\\ &=\kappa_{ez}\big((\mu_{2})_{xy}^{z}(\mu_{2})_{pq}^{y}\big)(\kappa^{-1})^{ax}(\kappa^{-1})^{bp}(\kappa^{-1})^{qc}s^{e}\,.\end{split}$
As in the previous case, this vanishes due to the fact that $\mu_{2}$ descends to a derivation on $V$, as shown in Proposition B.0.1. ∎
Now denote by ${\mathcal{K}}_{u}$ the raviolo polynomials (4.3.12) locally around the holomorphic variable $u=0$, such that its graded compoents are given by ${\mathcal{K}}_{u}^{0}={\mathbb{C}}[[u]]$ and ${\mathcal{K}}_{u}^{1}=\mathrm{span}_{\mathbb{C}}\{\Omega^{0}_{u},\Omega^{1}_{u},\dots\}$. With this at hand, we define the higher affine graded Lie algebra $\widehat{\mathfrak{G}}$ as
(6.1.9) $\widehat{\mathfrak{G}}=\left(V^{\vee}\otimes{\mathcal{K}}_{u}^{0}\right)\oplus\left(\mathfrak{n}^{\vee}\otimes{\mathcal{K}}_{u}^{1}\right).$
The bracket on $\widehat{\mathfrak{G}}$ is defined as follows: For $X\otimes f(u),Y\otimes g(u)\in V^{\vee}\otimes{\mathcal{K}}_{u}^{0}$, we set
(6.1.10) $X\otimes f(u),Y\otimes g(u)]=[X,Y]^{\vee}\otimes f(u)g(u)\in V^{\vee}\otimes{\mathcal{K}}_{u}^{0}\,.$
For $X\otimes f(u)\in V^{\vee}\otimes{\mathcal{K}}_{u}^{0}$ and $A\otimes k(\Omega_{u})\in\mathfrak{n}^{\vee}\otimes{\mathcal{K}}_{u}^{1}$, the bracket is
(6.1.11) $X\otimes f(u),A\otimes k(\Omega_{u})]=\mu_{2}^{\vee}(X,A)\otimes f(u)k(\Omega_{u})\in\mathfrak{n}^{\vee}\otimes{\mathcal{K}}_{u}^{1}\,.$
Finally, for two elements $A\otimes k(\Omega_{u}),B\otimes l(\Omega_{u})\in\mathfrak{n}^{\vee}\otimes{\mathcal{K}}_{u}^{1}$, the bracket vanishes:
(6.1.12) $A\otimes k(\Omega_{u}),B\otimes l(\Omega_{u})]=0\,.$
This bracket structure is consistent with the relations
(6.1.13) $u^{n}\Omega^{m}_{u}=\begin{cases}0&\text{if }n>m\,,\\ \Omega^{m-n}_{u}&\text{if }n\leq m\,,\end{cases}\qquad\Omega^{n}_{u}\Omega^{m}_{u}=0\,.$
With this definition at hand, we can show that the graded Lie brackets reproduce the commutators of the higher currents, up to the central term. Specifically, consider the bijective graded linear map
(6.1.14) ${\tilde{H}}_{n}^{a}\mapsto t^{a}\otimes u^{n}\,,\qquad{\tilde{L}}_{m}^{a}\mapsto s^{a}\otimes\Omega^{m}_{u}\,$,
which identifies the modes of the currents with the generators of the affine graded Lie algebra. Using the relations in equation (6.1.13), one readily verifies that the non-central contributions to the commutators in equations (5.2.7) and (5.2.8) coincide precisely with the graded Lie bracket on $\widehat{\mathfrak{G}}$.
6.2. The Central Extension
The final ingredient in the construction is the central extension. In the case of the Kac–Moody algebra, the central term arises from the $2$-cocycle (6.0.3), built from the invariant bilinear form on $\mathfrak{g}$ and the residue pairing defined by integration over $S^{1}$. In our setting, a similar construction yields a central extension of the affine graded Lie algebra $\widehat{\mathfrak{G}}$. To define the corresponding $2$-cocycle, we must first introduce the appropriate invariant bilinear form and an analogue of the residue pairing.
For the residue pairing, we will use the integration along $S^{2}$ introduced in the previous sections which satisfies
(6.2.1) $\frac{1}{8\pi i}\oint_{S^{2}}\mathrm{d}z\wedge u^{n}\Omega^{m}_{u}=\delta_{n,m}\,.$
With this, we can define a residue pairing as a map
$\operatorname{Res}=\operatorname{Res}_{u=0}:\mathcal{K}_{u}^{\bullet}\to\mathbb{C}[-1]$
which acts on the raviolo polynomial generators as
(6.2.2) $\operatorname{Res}(u^{n})=0\,,\quad\operatorname{Res}(u^{n}\Omega_{u}^{m})=\frac{1}{8\pi i}\oint_{S^{2}}\mathrm{d}u\wedge u^{n}\Omega^{m}_{u}=\delta_{n,m}\,.$
The notation ${\mathbb{C}}[-1]$ reflects the fact that we are working with chain complexes. Concretely, it indicates that $\operatorname{Res}$ vanishes on $(0,0)$-forms, while its action on $(1,0)$-forms yields values in $\mathbb{C}$}, as explicitly shown in (6.2.2).
On the other hand, to define a bilinear form on $\mathfrak{G}^{\vee}=\mathfrak{n}^{\vee}\oplus V^{\vee}$, we make use of the bilinear pairing on $H^{\bullet}(\mathfrak{G})=V\oplus\mathfrak{n}$, which we establish to be non-degenerate and invariant in Proposition B.0.1. Recall that, for a basis $\{t_{a}\}$ of $\mathfrak{n}$ and $\{s_{b}\}$ of $V$, this pairing is defined by $\langle t_{a},s_{b}\rangle=\kappa_{ab}$ where $\kappa_{ab}$ is an invertible matrix due to the non-degeneracy of the form. Taking the basis $\{s^{a}\}$ of $\mathfrak{n}^{\vee}$ and $\{t^{b}\}$ of $V^{\vee}$ we define the bilinear form $\langle\cdot,\cdot\rangle^{\vee}:V^{\vee}\times\mathfrak{n}^{\vee}\to{\mathbb{C}}$ by
(6.2.3) $\langle t^{a},s^{b}\rangle^{\vee}=\big(\kappa^{-1}\big)^{ab}\,.$
In the Kac–Moody case, the fact that $\omega$ defines a 2-cocycle relies crucially on the invariance of the bilinear form on $\mathfrak{g}$. The same principle applies here, so we now proceed to show that our bilinear pairing satisfies the required invariance condition.
6.2.1 Proposition.
The pairing $\langle\cdot,\cdot\rangle^{\vee}:V^{\vee}\times\mathfrak{n}^{\vee}\to{\mathbb{C}}$ is invariant.
Proof.
We aim to show that the bilinear pairing $\langle\cdot,\cdot\rangle^{\vee}$ on $\widehat{\mathfrak{G}}$ is invariant under the dual Lie bracket structure, specifically that
(6.2.4) $\langle[t^{a},t^{b}]^{\vee},s^{c}\rangle^{\vee}=\langle t^{a},\mu_{2}^{\vee}(t^{b},s^{c})\rangle^{\vee}\,.$
Using the explicit expressions for the dual brackets from (6.1.4) and the dual bilinear form from (6.2.3), this identity becomes
(6.2.5) $(\kappa^{-1})^{dc}\,\tilde{f}^{ab}_{d}=(\kappa^{-1})^{ad}\,(\tilde{\mu}_{2})^{bc}_{d}\,.$
Substituting the expressions for the dual structure constants in terms of the original ones as in (6.1.5), we find the condition
(6.2.6) $(\kappa^{-1})^{ax}(\kappa^{-1})^{by}f_{xy}^{c}=(\kappa^{-1})^{bx}(\mu_{2})^{a}_{xy}(\kappa^{-1})^{yc}\,.$
Now, recall that the invariance of the original bilinear pairing $\langle\cdot,\cdot\rangle:\mathfrak{n}\times V\to\mathbb{C}$ implies
(6.2.7) $\langle[t_{a},t_{b}],s_{c}\rangle=\langle t_{a},\mu_{2}(t_{b},s_{c})\rangle\,.$
Writing this in components we obtain
(6.2.8) $f_{ab}^{d}\,\kappa_{dc}=\kappa_{ad}\,(\mu_{2})_{bc}^{d}\,$,
which is precisely equivalent to (6.2.6) upon multiplying both sides by the inverse of $\kappa$. This completes the proof.
With both the residue pairing and the bilinear form in place, we are now ready to define the $2$-cocycle that determines the central extension of the affine graded Lie algebra. We consider, for $X\otimes f(u)\in V^{\vee}\otimes{\mathcal{K}}_{u}^{0}$ and $A\otimes k(\Omega_{u})\in\mathfrak{n}^{\vee}\otimes{\mathcal{K}}_{u}^{1}$
(6.2.9) $\mathfrak{w}(X\otimes f(u),A\otimes k(\Omega_{u}))=\langle X,A\rangle^{\vee}\,\operatorname{Res}\left(\partial_{u}f(u)\,k(\Omega_{u})\right),$
and extend it to be zero on pairs of elements both in $V^{\vee}\otimes{\mathcal{K}}_{u}^{0}$ and in $\mathfrak{n}^{\vee}\otimes{\mathcal{K}}_{u}^{1}$. Let us show that $\mathfrak{w}$ is indeed a $2$-cocycle
6.2.2 Proposition.
The bilinear map $\mathfrak{w}:\widehat{\mathfrak{G}}\times\widehat{\mathfrak{G}}\to{\mathbb{C}}$ determines a graded Lie algebra 2-cocycle.
Proof.
To verify that $\mathfrak{w}$ defines a graded Lie algebra $2$-cocycle, we must check that
(6.2.10) $\mathfrak{w}\big([t^{a}\otimes u^{n},\,t^{b}\otimes u^{m}]^{\vee},\,s^{c}\otimes\Omega^{k}_{u}\big)+\mathfrak{w}\big(\mu_{2}^{\vee}(s^{c}\otimes\Omega_{u}^{k},\,t^{a}\otimes u^{n}),\,t^{b}\otimes u^{m}\big)\\ +\mathfrak{w}\big(\mu_{2}^{\vee}(t^{b}\otimes u^{m},\,s^{c}\otimes\Omega_{u}^{k}),\,t^{a}\otimes u^{n}\big)=0\,.$
Using the structure constants and the identification $z^{n}\Omega^{m}=\Omega^{m-n}$, this becomes
(6.2.11) $\mathfrak{w}\big(\tilde{f}^{ab}_{d}\,t^{d}\otimes u^{n+m},\,s^{c}\otimes\Omega_{u}^{k}\big)+\mathfrak{w}\big((\tilde{\mu}_{2})^{ca}_{d}\,s^{d}\otimes\Omega_{u}^{k-n},\,t^{b}\otimes u^{m}\big)\\ +\mathfrak{w}\big((\tilde{\mu}_{2})^{bc}_{d}\,s^{d}\otimes\Omega^{k-m}_{u},\,t^{a}\otimes u^{n}\big)\,.$
Applying the definition of $\mathfrak{w}$ and the identity $\partial_{u}\Omega^{m}_{u}=-(m+1)\Omega_{u}^{m+1}$, we find
(6.2.12) $\tilde{f}^{ab}_{d}\,(\kappa^{-1})^{dc}(n+m)\,\delta_{n+m-1,k}-(\tilde{\mu}_{2})^{ca}_{d}\,(\kappa^{-1})^{bd}(k-n+1)\,\delta_{m,k-n+1}\\ -(\tilde{\mu}_{2})^{bc}_{d}\,(\kappa^{-1})^{da}(k-m+1)\,\delta_{n,k-m+1}\,.$
Next, we use the invariance of the bilinear form $\langle\cdot,\cdot\rangle^{\vee}$, which implies
(6.2.13) $\tilde{f}^{ab}_{d}\,(\kappa^{-1})^{dc}=(\tilde{\mu}_{2})^{bc}_{d}\,(\kappa^{-1})^{da}=(\tilde{\mu}_{2})^{ca}_{d}\,(\kappa^{-1})^{bd}\,.$
Substituting this into the previous expression, we obtain
(6.2.14) $\tilde{f}^{ab}_{d}\,(\kappa^{-1})^{dc}\left[(n+m)\,\delta_{n+m-1,k}-(k-n+1)\,\delta_{m,k-n+1}-(k-m+1)\,\delta_{n,k-m+1}\right]\\ =2\,\tilde{f}^{ab}_{d}\,(\kappa^{-1})^{dc}\,(n+m-k-1)\,\delta_{n+m-k-1,0}=0\,$,
which proves the 2-cocycle condition. ∎
It is straightforward to verify that this $2$-cocycle precisely reproduces the central term that appears in the commutator brackets of the higher currents:
(6.2.15) $\mathfrak{w}(t^{a}\otimes u^{n},s^{b}\otimes\Omega_{u}^{m})=\langle t^{a},s^{b}\rangle^{\vee}\mathrm{Res}(nu^{n-1}\Omega_{u}^{m})=n\big(\kappa^{-1}\big)^{ab}\delta_{n-1,m}\,.$
Using the identification (6.1.14), we see that this expression precisely matches the central term in the commutator given in (5.2.8). We have thus shown
6.2.3 Theorem.
The Poisson algebra of the currents of the 3d theory (4.2.1) defines a centrally extended affine Lie algebra $\widehat{\mathfrak{G}}_{\mathfrak{w}}$ fitting into the central extension sequence
(6.2.16) $0\rightarrow{\mathbb{C}}[1]\rightarrow\widehat{\mathfrak{G}}_{\mathfrak{w}}\rightarrow\widehat{\mathfrak{G}}\rightarrow 0\,$,
where the graded Lie bracket is given defined as follows: For $X\otimes f(u),Y\otimes g(u)\in V^{\vee}\otimes{\mathcal{K}}_{u}^{0}$, we have
(6.2.17) $X\otimes f(u),Y\otimes g(u)]=[X,Y]^{\vee}\otimes f(u)g(u)\,.$
For $X\otimes f(u)\in V^{\vee}\otimes{\mathcal{K}}_{u}^{0}$ and $A\otimes k(\Omega_{u})\in\mathfrak{n}^{\vee}\otimes{\mathcal{K}}_{u}^{1}$, the bracket is
(6.2.18) $X\otimes f(u),A\otimes k(\Omega_{u})]=\mu_{2}^{\vee}(X,A)\otimes f(u)k(\Omega_{u})+\langle X,A\rangle^{\vee}\mathrm{Res}(\partial_{u}f(u)k(\Omega_{u}))\,.$
Finally, for two elements $A\otimes k(\Omega_{u}),B\otimes l(\Omega_{u})\in\mathfrak{n}^{\vee}\otimes{\mathcal{K}}_{u}^{1}$, the bracket vanishes:
(6.2.19) $A\otimes k(\Omega_{u}),B\otimes l(\Omega_{u})]=0\,.$
Proof.
The proof follows from the bijective identification
(6.2.20) ${\tilde{H}}_{n}^{a}\mapsto t^{a}\otimes u^{n}\,,\quad{\tilde{L}}_{m}^{a}\mapsto s^{a}\otimes\Omega^{m}_{u}\,.$
7. The Raviolo Vertex Algebra
As discussed in the previous sections, much of the formalism developed there mirrors that of two-dimensional conformal field theories: the operator product expansion, radial quantisation, and the emergence of an infinite-dimensional symmetry algebra. A central feature in this setting, is the state-operator correspondence, which establishes a one-to-one relation between local operators in the field theory and states in the associated Hilbert space. It is this correspondence that ultimately allows one to pass consistently from the field-theoretic description to the Hilbert-space formulation.
In two dimensions, this formalism is rigorously captured by the notion of a vertex algebra. A vertex algebra provides a unified and mathematically precise framework that encompasses the various cases of interest—such as Virasoro and affine Kac–Moody algebras—while formalising many of the manipulations commonly used in practice. In particular, it offers a systematic way to organise and justify the procedures involved in the computation of physical quantities.
A central contribution of [22] is the introduction of the raviolo vertex algebra, a three-dimensional analogue of the standard vertex algebra. From the perspective of quantum field theory, it encodes the algebra of local operators in a theory that is partially holomorphic and partially topological, just as ordinary vertex algebras capture the operator algebra of two-dimensional chiral theories.
In this section we show how the currents $\tilde{L}$} and $\tilde{H}$}, together with their OPEs given in Theorem 5.1.2, can be used to construct a raviolo vertex algebra. For completeness, we begin by recalling the relevant definitions and results, and then proceed to the explicit construction of the raviolo vertex algebra associated with our three-dimensional theory (4.2.1).
7.1. A brief review of raviolo vertex algebras
Before turning to the structures that define a raviolo vertex algebra, we introduce the raviolo delta function, the analogue of the usual delta function on $\mathbb{C}$} but adapted to the geometry of $\mathbb{R}$$\times{\mathbb{C}}$. It is given by [22]
(7.1.1) $\Delta(z-w)=\sum_{n\geq 0}w^{n}\Omega_{z}^{n}-\sum_{n\geq 0}z^{n}\Omega_{w}^{n}\,$,
and satisfies
(7.1.2) $\frac{1}{8\pi i}\oint_{S^{2}}\mathrm{d}z\wedge\Delta(z-w)f(z)=f(w)$
for any $f\in\mathcal{K}_{\mathrm{poly}}^{\bullet}$. Further properties of this distribution can be found in [22]; we will only use its definition here.
With the delta distribution at hand, we consider a $\mathbb{Z}$-graded vector space $\mathcal{V}$, which will be the space of states of the theory. The grading is taken to be by cohomological degree. The essential object in this construction is the raviolo field which will allow us to organize the action of a given local operator on $\mathcal{V}$}.
Definition 7.1.1.
A raviolo field on $\mathcal{V}$ is an element
(7.1.3) $A(z)=\sum_{m\geq 0}z^{m}A^{+}_{m}+\Omega_{z}^{m}A^{-}_{m}\,$,
such that for any $v\in V$ there exist $N$ sufficiently large with
(7.1.4) $A^{-}_{m}v=0\,\quad\text{for all }\,m\geq N\,.$
We say that a field $A(z)$ is homogeneous of degree $|A|$ if, for every $m$, the mode $A^{+}_{m}:{\mathcal{V}}\to{\mathcal{V}}$ is an endomorphism of degree $|A|$, while $A^{-}_{m}:{\mathcal{V}}\to{\mathcal{V}}$ is an endomorphism of degree |A|1|A|-1. We denote the space of raviolo fields by $\mathcal{F}_{rav}({\mathcal{V}})$.
In a QFT, one is often interested in quantities such as $\langle\Psi|A(z)^{2}|\Psi\rangle$, the expectation value of the square of a quantum field $A(z)$ at a point $z$ in a state $|\Psi\rangle$. As seen explicitly in the OPE computations of the previous section, such expressions typically diverge. Physically, this reflects the fact that the fluctuations of $A(z)$ in the chosen state are infinite; equivalently, the variance of individual measurements of $A(z)$ is unbounded. What one ultimately seeks, however, is an operator whose classical limit is $A(z)^{2}$. To construct such an object, one must remove the short-distance divergences and retain a finite part, which can be interpreted as capturing the large-scale behaviour of the classical observable $A(z)^{2}$, rather than the short-scale fluctuations of the quantum field [43].
In this setting, the operator product expansion provides a precise description of the singular behaviour that appears when fields are multiplied at coincident points. This in turn motivates the definition of the normal ordered product, obtained by discarding the singular contributions.
Definition 7.1.2.
Given $A(z)$ and $B(z)$ raviolo fields on $\mathcal{V}$, we define their normal ordered product :A(z)B(w)::A(z)B(w): as
(7.1.5) $A(z)B(w):=A(z)_{+}B(w)+(-1)^{|A||B|}B(w)A(z)_{-}$
where $A_{+}(z)=\sum_{m\geq 0}z^{m}A^{+}_{m}$ and $A_{-}(z)=\sum_{m\geq 0}A^{-}_{m}\Omega^{m}$
The definition of the normal ordered product is entirely analogous to the two-dimensional case. It is given by the operator product expansion with the singular terms removed, that is, by the regular part of the OPE.
Normal ordering gives a consistent prescription for defining composite operators at coincident points. The other key structural ingredient is locality. In quantum field theory, locality means that the commutator of two fields vanishes when evaluated at distinct spacetime points, so that operators at spacelike separation commute. In the present setting, as seen in the OPEs computed in section §5, the number of singular terms is always finite. This reflects the strong constraints imposed by the partial holomorphic–topological symmetry, which tightly restricts the possible short-distance behaviour of composite operators. It is precisely this rigidity that makes it possible to define a consistent algebraic structure governing the product of fields—something that would not be available in a more general quantum field theory. More precisely, we have
Definition 7.1.3.
Given raviolo fields $A(z)$ and $B(w)$ on $\mathcal{V}$ we say that they are mutually local if there exist $N\geq 0$ such that
(7.1.6) $(z-w)^{N+1}[A(z),B(w)]=0$
and
(7.1.7) $\left(\Omega_{z}^{m}-\sum_{n=0}^{N}(w-z)^{n}\binom{m+n}{n}\Omega_{w}^{m+n}\right)[A(z),B(w)]=0\,.$
As shown in [22], mutual locality admits several equivalent formulations, closely mirroring those found in the theory of vertex algebras [44, 45]. The above definition makes it explicit that mutually local fields possess an OPE with only finitely many singular terms. In practice, however, it will be more convenient to use an equivalent formulation, stated in the following lemma.
7.1.4 Lemma.
The raviolo fields $A(z)$ and $B(w)$ are mutually local if and only if there exist raviolo fields $C_{n}(w)$ such that
(7.1.8) $A(z),B(w)]=\sum_{n=0}^{N}\frac{1}{n!}\partial_{w}^{n}\Delta(z-w)C_{n}(w)$
A proof of this lemma can be find in Proposition 2.2.2 of [22].
Together, normal ordering and mutual locality ensure that the collection of fields is closed under products: composite operators can be defined consistently, and their short-distance singularities are always under control. Beyond this, and in close analogy with the two-dimensional case, one can establish a one-to-one identification between raviolo fields and states in the vector space $\mathcal{V}$. Thus, in addition to possessing an algebra of operators, the theory is equipped with a vector space structure that organises these operators in a systematic way. This feature is known as the state–operator correspondence, realised as a linear map from $\mathcal{V}$ to the space of raviolo fields on $\mathcal{V}$. That is,
(7.1.9) $Y(-,z):{\mathcal{V}}\to{\mathcal{F}}_{rav}({\mathcal{V}})\,$,
such that for every $a\in{\mathcal{V}}$ we have
(7.1.10) $Y(a,z)=\sum_{m\geq 0}z^{m}a^{+}_{m}+\sum_{m\geq 0}\Omega_{z}^{m}a^{-}_{m}\,.$
With all of these definitions at hand, we can introduce the main definition of a raviolo vertex algebra following [22]
Definition 7.1.5.
A raviolo vertex algebra is the data $(\mathcal{V},|0\rangle,\partial,Y)$ where
  • ${\mathcal{V}}=\oplus{\mathcal{V}}^{r}$ is the space of states, which is $\mathbf{Z}$}-graded
  • $|0\rangle$ is the vacuum vector which is a distinguished element in ${\mathcal{V}}^{0}$
  • $\partial:{\mathcal{V}}\to{\mathcal{V}}$ is the translator operator, given by a degree 0 endomorphism
  • $Y:{\mathcal{V}}\to\mathcal{F}_{rav}({\mathcal{V}})$ is a linear map of degree 0
subject to the following axioms
  1. (1)
    For every $a\in{\mathcal{V}}$ the element $Y(a,z)$ is a raviolo field
  2. (2)
    For every $a\in{\mathcal{V}}$ we have
    $\partial,Y(a,z)]=\partial_{z}Y(a,z)$
  3. (3)
    The vacuum vector satisfies $\partial|0\rangle=0$, $Y(|0,\rangle,z)=\mathbb{1}_{{\mathcal{V}}}$, and for every $a\in{\mathcal{V}}$ we have $Y(a,z)|0\rangle\in\mathcal{V}[\![z]\!]$; that is, $Y_{-}(a,z)|0\rangle=0$.
  4. (4)
    For every $a,b\in{\mathcal{V}}$ the fields $Y(a,z)$ and $Y(b,w)$ are mutually local.
The definition of a raviolo vertex algebra involves the graded vector space $\mathcal{V}$}, the state–operator correspondence, and the mutual locality of fields. These elements are sufficient to define the fields of the theory and to prescribe how they compose through their operator products. To capture the full physical structure of the theory, however, two further ingredients are required. The first is the vacuum state, representing the ground state of the theory, whose axioms combine the expected physical features of the vacuum with the conditions needed for mathematical consistency. The second is the translation operator $\partial$, which generates infinitesimal displacements along the complex direction. It allows one to translate the notion of differentiating fields into an algebraic operation on states.
With these ingredients in place, one obtains a complete algebraic description of the local operators of a three-dimensional theory with the symmetry described in the previous sections. Much as in the two-dimensional case of vertex algebras, this structure makes it possible to relate representation theory to physical quantities, thereby extending many of the familiar applications of chiral vertex algebras to the raviolo setting.
In practice, the existence of a raviolo vertex algebra is most easily established by means of a reconstruction theorem, in direct analogy with the two-dimensional case [46]. Such theorems provide an alternative set of conditions which guarantee the existence of a raviolo vertex algebra structure, but are often simpler to verify in concrete situations. More precisely, following [22] we have
7.1.6 Proposition.
Let $\mathcal{V}=\bigoplus\mathcal{V}^{r}$ be a $\mathbb{Z}$–graded vector space, $|0\rangle$ a non–zero vector, $\partial$ a degree 0 endomorphism of $\mathcal{V}$. Further, let $\{a^{i}\}$ be a countable ordered set of vectors in $\mathcal{V}$, with $a^{i}\in\mathcal{V}^{r_{i}}$, and suppose we are given homogeneous fields
(7.1.11) $A^{i}(z)=\sum_{m\geq 0}z^{m}A^{i,+}_{m}+\Omega^{m}_{z}A^{i,-}_{m}$
of degree $r_{i}$, such that the following hold:
  1. (1)
    $A^{i,+}_{0}|0\rangle=a^{i}$ and $A^{i,-}_{m}|0\rangle=0$ for all $i$ and $m\geq 0$.
  2. (2)
    $\partial|0\rangle=0$ and $\partial,A^{i}(z)]=\partial_{z}A^{i}(z)$ for all $i$.
  3. (3)
    All fields $A^{i}(z)$ are mutually local.
  4. (4)
    $\mathcal{V}$ is spanned by the vectors
    (7.1.12) $A^{i_{1}}_{j_{1}}\cdots A^{i_{l}}_{j_{l}}|0\rangle,\qquad j_{k}\geq 0\,.$
Then the assignment
(7.1.13) $Y\!\left(A^{i_{1}}_{j_{1}}\cdots A^{i_{l}}_{j_{l}}|0\rangle,z\right)\\ =\frac{1}{(j_{1}-1)!\cdots(j_{l}-1)!}:\partial_{z}^{-j_{1}-1}A^{i_{1}}(z)\cdots\partial_{z}^{-j_{l}-1}A^{i_{l}}(z):$
determines a well-defined raviolo vertex algebra structure on $\mathcal{V}$. Moreover, this is the unique raviolo vertex algebra structure on $\mathcal{V}$ satisfying conditions (1)–(4) and such that $Y(a^{i},z)=A^{i}(z)$.
7.2. Raviolo Current Vertex Algebra
In this final section, we construct the raviolo current vertex algebra associated with the anti-chiral sector of the topological-holomorphic theory (4.2.1). Our strategy is to leverage the canonical representation of $\widehat{\mathfrak{G}}_{\mathfrak{w}}$ on its universal envelope, then apply the reconstruction theorem (Prop. 4.0.1 of [22]) introduced above.
Let us begin by constructing the vacuum module $\mathcal{V}$. From a physicist’s perspective, the vacuum module is simply the Fock space — the vector space generated by arbitrary products of creation operators acting on the vacuum. Since our raviolo vertex algebra will be built from the affine graded Lie algebra $\widehat{\mathfrak{G}}_{\mathfrak{w}}$ constructed in §6, we must identify, among its generators, which act as creation and which as annihilation operators.
From point (1) of Proposition 7.1.6, the creation operators are the $A^{i,+}_{n}$’s, which in the raviolo field expansion come proportional to powers of $z$. The annihilation operators, on the other hand, are the $A^{i,-}_{n}$’s which come proportional to the form-valued basis elements $\Omega_{z}^{n}$. Given the raviolo fields are homogeneous in cohomological degree, their components must pair consistently with $z$ and $\Omega_{z}$. This determines that the generators $\tilde{L}_{n}^{a}$ are paired with $z$ and thus act as creation operators, while the generators $\tilde{H}_{n}^{a}$ are paired with $\Omega_{z}$ and act as annihilation operators.
The vacuum module $\mathcal{V}$ is therefore spanned by states of the form
(7.2.1) $\tilde{L}_{j_{1}}^{a^{1}}\tilde{L}_{j_{2}}^{a_{2}}\cdots\tilde{L}_{j_{k}}^{a_{k}}|0\rangle\,,\qquad j_{i}\geq 0,$
and such that
(7.2.2) $\tilde{H}_{n}^{a}|0\rangle=0\,,\qquad n\geq 0\,.$
To formalize this intuitive picture, one can describe the construction of $\mathcal{V}$ in purely algebraic terms. Consider the affine graded Lie algebra $\widehat{\mathfrak{G}}_{\mathfrak{w}}$ generated by the elements $\ell^{a}_{n}=s^{a}\otimes\Omega_{u}^{n}$ in degree 1 and $\eta^{a}_{n}=t^{a}\otimes u^{n}$ in degree 0. Its universal enveloping graded algebra, denoted $U\widehat{\mathfrak{G}}_{\mathfrak{w}}$, is defined as the free graded algebra generated by polynomials in $\ell_{n}^{a}$ and $\eta_{n}^{a}$, together with a central element $\mathfrak{w}$ of degree 1.
The generators are ordered according to the graded Poincaré–Birkhoff–Witt theorem, and obey the affine Lie algebra relations
$\ell^{a}_{n}\ell^{b}_{m}-\ell^{b}_{m}\ell^{a}_{n}=0\,,\qquad\eta^{a}_{n}\eta^{b}_{m}-\eta^{b}_{m}\eta^{a}_{n}=[\eta^{a}_{n},\eta^{b}_{m}]\,$,
$\eta^{a}_{n}\ell^{b}_{m}-\ell^{b}_{m}\eta^{a}_{n}=[\eta^{a}_{n},\ell^{b}_{m}]\,.$
The homogeneous degree-0 unit is $(\ell^{a}_{n})^{0}=(\eta^{a}_{m})^{0}=1$. By working under the identification (6.2.20),
(7.2.3) ${\tilde{L}}_{n}^{a}\mapsto\ell^{a}_{n}=s^{a}\otimes\Omega_{u}^{n}\,,\quad{\tilde{H}}_{n}^{a}\mapsto\eta^{a}_{n}=t^{a}\otimes u^{n}$
the mode generators ${\tilde{L}}^{a}_{n},{\tilde{H}}^{a}_{n}$ act canonically on $U\hat{\mathfrak{G}}_{\mathfrak{w}}$ by multiplication by $\ell^{a}_{n},\eta^{a}_{n}$, respectively. We will use (7.2.3) extensively in the following.
Next, consider the central factor $\mathbb{C}_{w}=\mathbb{C}\oplus\mathbb{C}w[-1]$, where $w$ is an abstract Abelian generator of degree 1, and whose homogeneous degree-0 unit is $1\oplus 0w=1$. Taking the tensor product $U\widehat{\mathfrak{G}}_{\mathfrak{w}}\otimes\mathbb{C}_{w}$, the vacuum module $\mathcal{V}$ is obtained by quotienting out two relations: (i) all of the $\tilde{H}$}-modes act (through $\rho$) trivially on $\mathcal{V}$, and (ii) the generator of the central piece $\mathbb{C}[1]\subset\widehat{\mathfrak{G}}_{\mathfrak{w}}$ acts as multiplication by the constant ww. [^10] Note that both the values of the 2-cocycle $\mathfrak{w}$ and the scalar $w$ have cohomological degree 1. Equivalently,
(7.2.4) $\mathcal{V}=U\widehat{\mathfrak{G}}_{\mathfrak{w}}\otimes_{U\widehat{\mathfrak{G}}_{\geq 0}}\mathbb{C}_{w}\,$,
where $\widehat{\mathfrak{G}}_{\geq 0}=\mathrm{span}\{{\tilde{H}}_{n}^{a}\}\oplus\mathbb{C}w[-1]$ denotes the positive graded subalgebra.
The following equivalent mathematical definition should be familiar to experts in vertex algebras.
Definition 7.2.1.
The vacuum module $\mathcal{V}=\operatorname{Ind}_{U\hat{\mathfrak{G}}_{\geq 0}}^{U\hat{\mathfrak{G}}_{\mathfrak{w}}}{\mathbb{C}}_{w}$ is the induced representation of the trivial $\hat{\mathfrak{G}}_{\mathfrak{w}}$-representation on ${\mathbb{C}}_{w}\cong{\mathbb{C}}\oplus{\mathbb{C}}[1]$.
In the following, we shall leverage (7.2.3) in order to distinguish a vector $\ell^{a}_{n},\eta^{a}_{n}\in\widehat{\mathfrak{G}}_{\mathfrak{w}}\subset\mathcal{V}$, and its corresponding operator ${\tilde{L}}^{a}_{n},{\tilde{H}}^{a}_{n}$ on $\mathcal{V}$.
We are now in conditions to prove the following
7.2.2 Theorem.
The conserved currents arising from the three-dimensional action (4.2.1) give rise to a raviolo vertex algebra structure.
Proof.
Let us take the vector space $\mathcal{V}$ defined in (7.2.4) and consider, for $1\leq a\leq\operatorname{dim}\mathfrak{n}=\operatorname{dim}V$, the raviolo field operators
(7.2.5) $\mathcal{J}^{a}(z)=\sum_{n\geq 0}z^{n}{\tilde{L}}_{n}^{a}+\Omega_{z}^{n}{\tilde{H}}^{a}_{n}\,$,
which are homogeneous of degree 0. We identify the unit $|0\rangle=1\otimes 1\in\mathcal{V}$ as the vacuum. We will now prove points (1) through (4) of Proposition 7.1.6.
By definition of $\mathcal{V}$, we have that
(7.2.6) ${\tilde{H}}_{n}^{a}|0\rangle=0\,,\quad{\tilde{L}}_{0}^{a}|0\rangle=\ell_{0}^{a}\,$,
which proves point (1). Next, for the translation operator, we begin by defining an additional generator $D$ on $\widehat{\mathfrak{G}}_{\mathfrak{w}}\cong(\mathfrak{G}^{\vee}\otimes\mathcal{K}^{\bullet}_{u})\oplus{\mathbb{C}}[1]$ from the action of the derivative $\partial_{u}$ on the current modes:
$D,X\otimes k(u)]=-X\otimes\partial_{u}k(u),\qquad[D,\mathfrak{w}]=0\,$,
where $X\in\mathfrak{G}^{\vee},~k(u)\in\mathcal{K}_{u}^{\bullet}$ and $\mathfrak{w}\in{\mathbb{C}}[1]$ is the generator. Under the canonical $\widehat{\mathfrak{G}}_{\mathfrak{w}}$-representation on $\mathcal{V}$, this gives rise to an endomorphism $\partial\in\operatorname{End}\mathcal{V}$ which, by construction, satisfies
(7.2.7) $\partial,x]=\partial_{z}x\,,\qquad\partial|0\rangle=0$
for all $x\in\widehat{\mathfrak{G}}_{\mathfrak{w}}$. More explicitly, these read as the following mode relations
(7.2.8) $\partial,{\tilde{L}}_{n}^{a}]=(n+1){\tilde{L}}_{n+1}^{a}\,,\quad[\partial,{\tilde{H}}_{n}^{a}]=-n{\tilde{H}}_{n-1}^{a}\,.$
This then allows us to compute
[…]
identifying $\partial$ as the desired translation operator, thereby proving point (2).
By performing a detailed computation that properly incorporates the grading, we verify that the commutator between any two raviolo fields satisfies
(7.2.10) $\mathcal{J}^{a}(z),\mathcal{J}^{b}(w)]=\partial_{w}\Delta(z-w)(\kappa^{-1})^{ab}+\Delta(z-w)\tilde{\textbf{f}}^{ab}_{c}\mathcal{J}^{c}(w)\,$,
for some constants $\tilde{\textbf{f}}^{ab}_{c}$ and thus, by Lemma (7.1.4) we obtain their mutual locality thereby proving point (3).
Finally, point (4) follows inmmediately by construction of $\mathcal{V}$, and thus, by the reconstruction theorem 7.1.6, there exist a unique raviolo vertex algebra structure on $\mathcal{V}$.
The expression of (7.2.10) bears a striking resemblance to the commutator bracket obtained for the ”raviolo current algebra” in [22]. However, one crucial difference is that our pairing form $\kappa^{-1}$ has odd degree. This non-trivial degree within $\kappa$ is in fact closely related to solutions of the 2-graded classical Yang-Baxter equations [47, 48, 49], as well as the construction of higher-braiding structures [50, 51] — both aspects relating back to higher-integrability (cf. [40]).
The story is not over yet, however — the full quantum operator products of the theory $\mathcal{W}$} involves independent higher-order products coming from homotopy transfer (see footnote 8 on pg. 21, and §B). The algebraic structure of these higher-order OPEs in higher-dimensional (ie. $\geq 1+1$d) quantum field theories has been well-studied recently [52]. For holomorphic field theories, these are known as the ”operatope” [53, 54], and a model for the involved higher-order operator products has been given in [55].
Following the same philosophy, the complete higher-order vertex algebraic quantization of $\mathcal{W}$} and its quantum higher-integrability will be studied in a future work.
Acknowledgements
Part of the completion of this project was supported by the China international talents Exchange Program (CEP code JC202502007G) at the Beijing Institute of Mathematical Sciences and Applications. JL thanks Lewis Cole for valuable comments on the introduction, and Horacio Falomir for his support as PhD advisor. The work of JL is supported by CONICET. HC is supported by the National Science Foundation of China (grant number W2533012).
Appendix A Conventions
Throughout this article, we work extensively with various cohomology groups. To keep our conventions transparent and accessible, we summarize them here so that the reader can refer back as needed to verify definitions and computations.
We focus primarily on two complexes, the first of which is the Dolbeault complex on $\mathbb{C}\setminus\{0\}$, organized as follows:
(A.0.1) $\Omega^{(0,0)}({\mathbb{C}}\setminus\{0\})$$\Omega^{(1,0)}({\mathbb{C}}\setminus\{0\})$$\cdots$$\Omega^{(0,1)}({\mathbb{C}}\setminus\{0\})$$\Omega^{(1,1)}({\mathbb{C}}\setminus\{0\})$$\cdots$$\scriptstyle{\partial}$$\scriptstyle{\bar{\partial}}$$\scriptstyle{\bar{\partial}}\scriptstyle{\partial}$
In this context, we define the $\partial$-cohomology groups by
(A.0.2) $H_{\partial}^{(p,q)}(\mathbb{C}\setminus\{0\})=\frac{\{\text{$\partial$-closed }(p,q)\text{-forms}\}}{\{\text{$\partial$-exact }(p,q)\text{-forms}\}}\,$,
and analogously for the $\bar{\partial}$-cohomology, by replacing $\partial$ with $\bar{\partial}$ in the definition above.
The second complex we will frequently use is the raviolo complex. Given a three-manifold $M$ equipped with a transverse holomorphic foliation, the raviolo complex is defined as follows:
(A.0.3) $\mathcal{A}^{0,0}(M)$$\mathcal{A}^{1,0}(M)$$\cdots$$\mathcal{A}^{0,1}(M)$$\mathcal{A}^{1,1}(M)$$\cdots$$\scriptstyle{{\mathrm{d}}^{\prime}}\scriptstyle{\partial}\scriptstyle{\partial}$$\scriptstyle{{\mathrm{d}}^{\prime}}$
[…]
(A.0.4) $H_{{\mathrm{d}}^{\prime}}^{(p,q)}(M)=\frac{\{\text{${\mathrm{d}}^{\prime}$-closed }(p,q)\text{-forms}\}}{\{\text{${\mathrm{d}}^{\prime}$-exact }(p,q)\text{-forms}\}}\,\,$,
and analogously for the $\partial$-cohomology, by replacing $\mathrm{d}$$^{\prime}$ with $\partial$ in the definition above.
The same notational conventions apply to all other complexes appearing throughout the text.
Appendix B The Cohomology of $\mathfrak{G}$
In the present case, the currents and transformation parameters are valued in the cohomology $H^{\bullet}(\mathfrak{G})=V\oplus\mathfrak{n}$, where, recall,
(B.0.1) $V=\ker(\mu_{1})\subset\mathfrak{h}\,,\qquad\mathfrak{n}=\mathfrak{g}/\operatorname{im}(\mu_{1})\,.$
To write expressions like $\langle\tilde{\alpha},\tilde{H}\rangle$ or $\tilde{\alpha},\tilde{\alpha}^{\prime}$], as in the two-dimensional case, it is necessary to verify that these operations descend to cohomology. We include a proof of this fact for completeness; readers who are happy to take it as given may proceed directly to the Poisson bracket computation.
B.0.1 Proposition.
The following properties hold
  1. (1)
    The induced graded brackets on $V\oplus\mathfrak{n}$ satisfy the graded Jacobi identity.
  2. (2)
    The non-degenerate bilinear form $\langle\cdot,\cdot\rangle:\mathfrak{g}\times\mathfrak{h}\to{\mathbb{C}}$ induces a non-degenerate bilinear form $\langle\cdot,\cdot\rangle:\mathfrak{n}\times V\to{\mathbb{C}}$.
Note that the first property guarantees that $\mathfrak{n}$ inherits a Lie bracket from $\mathfrak{g}$, and that $\mu_{2}$ induces a well-defined derivation on $V$.
Proof.
There are natural projection and inclusion maps
$p=(p_{-1},p_{0}):\mathfrak{G}\to H^{\bullet}(\mathfrak{G}),\qquad\iota=(\iota_{-1},\iota_{0}):H^{\bullet}(\mathfrak{G})\to\mathfrak{G},$
which satisfy $p\circ\iota=\mathbb{1}_{H^{\bullet}(\mathfrak{G})}$. The equation $\iota\circ p\simeq\mathbb{1}_{\mathfrak{G}}$, on the other hand, only holds up to homotopy [56] (see also [57]): there exists a linear map $h:\mathfrak{g}\to\mathfrak{h}$, called the chain homotopy, which satisfies
(B.0.2) $h\circ\mu_{1}=\iota_{-1}\circ p_{-1}-\mathrm{id}_{\mathfrak{h}}\,,\qquad\mu_{1}\circ h=\iota_{0}\circ p_{0}-\mathrm{id}_{\mathfrak{g}}\,.$
This is the essence of the homotopy transfer theorem [58].
To prove (1) we can use this chain homotopy to show that the canonically induced Lie bracket on $\mathfrak{n}$ and the induced action on $V$
(B.0.3) $\cdot,\cdot]^{\prime}=p_{0}\circ[\cdot,\cdot]\circ(\iota_{0}\otimes\iota_{0})\qquad\mu_{2}^{\prime}=p_{-1}\circ\mu_{2}\circ(\iota_{0}\otimes\iota_{-1})$
satisfy the graded Jacobi identity.
Indeed, to see this, we compute for each $\mathsf{x},\mathsf{x}^{\prime}\in\mathfrak{n}$ and $\mathsf{y}\in V$ that
[…]
where we have used the identity $\mu_{1}\mu_{2}(\cdot,\cdot)=[\cdot,\mu_{1}\cdot]$, and the first term vanishes in the last line due to the definition of $\iota_{-1}:V=\operatorname{ker}(\mu_{1})\hookrightarrow\mathfrak{h}$.
Similarly, we compute for $\mathsf{x}_{1},\mathsf{x}_{2},\mathsf{x}_{3}\in\mathfrak{n}$,
[…]
where the first term vanishes in the last line due to the definition $p_{0}:\mathfrak{g}\rightarrow\mathfrak{n}=\mathfrak{g}/\operatorname{im}(\mu_{1})$. By summing over cyclic permutations, the remaining quantities in both of the above equations vanish by the graded Jacobi identity satisfied by $\mathfrak{G}$. Thus, property (1) holds.
To prove (2), we must show that the induced bilinear form $\langle\cdot,\cdot\rangle$ is both well defined and non-degenerate. First, we verify that it is independent of the choice of representatives. Let $\mathsf{x},\mathsf{x}^{\prime}\in\mathfrak{g}$ be representatives of the same class in $\mathfrak{n}=\mathfrak{g}/\mathrm{im}(\mu_{1})$, so that $\mathsf{x}^{\prime}=\mathsf{x}+\mu_{1}(\theta)$ for some $\theta\in\mathfrak{h}$. Then, for any $\mathsf{y}\in V=\ker(\mu_{1})$, we have
(B.0.4) $\langle\mathsf{x}^{\prime},\mathsf{y}\rangle=\langle\mathsf{x}+\mu_{1}(\theta),\mathsf{y}\rangle=\langle\mathsf{x},\mathsf{y}\rangle+\langle\mu_{1}(\theta),\mathsf{y}\rangle\,.$
The second term vanishes due to the compatibility of the bilinear form with $\mu_{1}$, namely,
(B.0.5) $\langle\mu_{1}(\theta),\mathsf{y}\rangle=\langle\theta,\mu_{1}(\mathsf{y})\rangle=0\,$,
since $\mathsf{y}\in\ker(\mu_{1})$. Therefore, the pairing is well defined on cohomology classes.
To show that the induced form is non-degenerate, we apply the rank-nullity theorem. Since $\mu_{1}:\mathfrak{h}\to\mathfrak{g}$ is linear, we have:
(B.0.6) $\dim\ker(\mu_{1})+\dim\mathrm{im}(\mu_{1})$ $=\dim\mathfrak{h}\,$,
(B.0.7) $\dim\mathrm{coker}(\mu_{1})+\dim\mathrm{im}(\mu_{1})$ $=\dim\mathfrak{g}\,.$
Since the original bilinear form is invertible we have that $\dim\mathfrak{h}=\dim\mathfrak{g}$ so that subtracting these equations gives
(B.0.8) $\dim V=\dim\ker(\mu_{1})=\dim\mathrm{coker}(\mu_{1})=\dim\mathfrak{n}\,$,
which implies that the restriction of $\langle\cdot,\cdot\rangle$ to $V\oplus\mathfrak{n}$ is non-degenerate.
For notational simplicity, we will use the same symbols $\cdot$$\cdot$], $\mu_{2}$, and $\langle\cdot,\cdot\rangle$ for the induced bracket, action, and bilinear form on $H^{\bullet}(\mathfrak{G})$, omitting primes to avoid clutter. Moreover, we will denote the corresponding structure constants using unprimed letters. Given a basis $\{t_{a}\}$ of $\mathfrak{n}$ and $\{s_{b}\}$ of $V$, we thus write:
(B.0.9) $t_{a},t_{b}]=f_{ab}^{c}\,t_{c}\,,\qquad\mu_{2}(t_{a},s_{b})=(\mu_{2})_{ab}^{c}\,s_{c}\,,\qquad\langle t_{a},s_{b}\rangle=\kappa_{ab}\,.$